### v1

# Active Fine-Tuning of Generalist Policies

Abstract
Pre-trained generalist policies are rapidly gaining relevance in robot learning due to their promise of fast adaptation to novel, in-domain tasks. This adaptation often relies on collecting new demonstrations for a specific task of interest and applying imitation learning algorithms, such as behavioral cloning. However, as soon as several tasks need to be learned, we must decide which tasks should be demonstrated and how often? We study this multi-task problem and explore an interactive framework in which the agent adaptively selects the tasks to be demonstrated. We propose AMF (Active Multi-task Fine-tuning), an algorithm to maximize multi-task policy performance under a limited demonstration budget by collecting demonstrations yielding the largest information gain on the expert policy. We derive performance guarantees for AMF under regularity assumptions and demonstrate its empirical effectiveness to efficiently fine-tune neural policies in complex and high-dimensional environments.
1 Introduction
Figure 1: Interactive loop between agent and expert. We consider a scenario where we receive a pre-trained policy, and are able to obtain expert demonstrations of tasks. We study how to select tasks (in blue) to obtain the best-performing policy after as few demonstrations as possible.
The availability of large pre-trained models has transformed entire areas of machine learning, from computer vision (Krizhevsky et al., 2012; He et al., 2016; Dosovitskiy et al., 2021; Radford et al., 2021), to natural language processing (Radford et al., 2019; Brown et al., 2020) and generative modeling in general (Ho et al., 2020; Esser et al., 2024). This paradigm has started to extend to robotics and control (Collaboration, 2023; Ma et al., 2024), in particular for systems for which demonstrations are readily available (Octo Model Team et al., 2024), or can be easily collected (Zhao et al., 2023). Even when demonstrations are not easily obtained, scaling laws in reinforcement learning (Ceron et al., 2024b; a; Nauman et al., 2024) suggest the possibility of leveraging large pre-trained policies. These “generalist” policies have decent performance on many tasks, and can be fine-tuned on particular set of tasks while leveraging their previously learned representations and skills. We investigate whether representations of such policies can be used to significantly bootstrap learning progress.
As a motivating example, consider a household robot that is delivered with a pre-trained “generalist” policy, and deployed in slightly different conditions than those observed in its training data. While the robot may achieve some tasks in a zero-shot fashion (e.g. simple pick-and-place), other tasks might necessitate further fine-tuning (e.g. cooking an omelette). The robot should be able to interactively request demonstrations to compensate for its shortcomings. We seek to answer which demonstrations should be requested from the user to achieve the best performance, as quickly as possible.
If the agent only needs to perform well in a single task, the fine-tuning process conventionally relies on behavioral cloning (Chen et al., 2021; Reed et al., 2022; Bousmalis et al., 2024) of expert demonstrations. As collecting demonstrations is in general costly, the number of demonstrations required, and thus the expert’s effort, should be minimized. However, as each demonstration should solve the same task, the allocation of the expert’s effort is straightforward. The multi-task case presents the more nuanced problem of selecting which tasks to demonstrate, and when. This motivates the main focus of this work: provided a pre-trained policy, how can we maximize multi-task performance with a minimal number of additional demonstrations?
To address this problem, we propose AMF (Active Multi-task Fine-tuning), which selects maximally informative demonstrations. AMF parallels recent work on active supervised fine-tuning of neural networks (Hübotter et al., 2024). To this end, AMF relies on estimates of the demonstrations’ information gain about the expert policy. We prove that in sufficiently regular Markov decision processes, AMF converges to the expert policy. We then focus on practical scenarios where policies are represented as neural networks. We show that, despite additional challenges, AMF can effectively guide active task selection in such settings, leading to better policies after fewer demonstrations. Our contributions are:
  • We propose AMF, an algorithm for multi-task policy fine-tuning that maximizes the information gain of demonstrations about the expert policy.
  • We prove statistical guarantees for AMF, which extend the results of Hübotter et al. (2024) to dynamical systems.
  • We empirically scale AMF to high-dimensional tasks involving pre-trained neural policies.
2 Related Work
Learning-based control and active data selection are both well-established research directions. This section discusses some of the most topical works in either direction, and clarifies the novelty and placement of this work with respect to them.
Behavioral Cloning
Numerous imitation learning approaches have been developed with the goal of distilling knowledge from high-quality demonstrations to a control policy (Osa et al., 2018). Within this family of techniques, behavioral cloning (BC, Bain & Sammut, 1995; Ross & Bagnell, 2010) aims to maximize policy performance by minimizing the distance of its actions to demonstrated actions, simply through supervised learning. While BC may suffer from accumulating errors (Ross et al., 2011), its empirical effectiveness has seen increasing support when high-quality demonstrations are readily available (Kumar et al., 2022). Next to recent empirical successes (Chi et al., 2023), formal analysis has also advanced (Spencer et al., 2021; Block et al., 2024a; Belkhale et al., 2024; Foster et al., 2024), and established provable performance guarantees for BC policies (Xu et al., 2020; Maran et al., 2023; Block et al., 2024b).
Multi-task and Generalist Policies
Traditionally, behavioral cloning has mostly been deployed in a single-task setting. Multi-task learning in sequential decision-making has largely been investigated in the context of reinforcement learning (Teh et al., 2017; Sodhani et al., 2021; Yu et al., 2021; Sun et al., 2022; Cho et al., 2022; Hendawy et al., 2023). Moreover, the recent rise of multi-task generative models (Brown et al., 2020) has been mirrored by exploration of multi-task, or generalist policies, often trained via imitation learning (Reed et al., 2022; Bousmalis et al., 2024; Collaboration, 2023). These recent works mostly build upon algorithms developed for the single-task case, and simply integrate task-conditioning as part of the state. While several works hand-select parts of large, open-source robotics datasets for pre-training (Octo Model Team et al., 2024), active data selection for multi-task fine-tuning has not been addressed. Prior work on meta-learning has studied how one can explicitly meta-learn the ability to adapt to task demonstrations (Finn et al., 2017). We find this capability to emerge even from models that are not explicitly trained in this way, and focus on which demonstrations to obtain.
Data Selection
The idea of directing a sampling process to gather information has been central to machine learning research and studied extensively in experimental design (Chaloner & Verdinelli, 1995) and active learning (Settles, 2009). Most work on active data selection summarizes data without focusing on a particular task (e.g., Sener & Savarese, 2017; Ash et al., 2020b; Holzmüller et al., 2023; Lightman et al., 2023), which has been predominantly applied to pre-training. Recently, adapting models after pre-training and during deployment has gained interest. Several works, mostly in computer vision, focus on unsupervised fine-tuning on a test instance (Jain & Learned-Miller, 2011; Krause et al., 2018; Sun et al., 2020; Wang et al., 2021b; Chen et al., 2022). We focus instead on supervised fine-tuning of learning-based controllers in dynamical systems. This necessitates automatic data selection, for which practical methods currently rely on uniform sampling or externally provided heuristics. Our approach extends work on task-directed data selection (Kothawade et al., 2020; Wang et al., 2021a; Kothawade et al., 2022; Smith et al., 2023), which has recently been applied to the supervised fine-tuning of large-scale neural networks in vision (Hübotter et al., 2024) and language (Xia et al., 2024).
3 Background
3.1 Multi-Task Reinforcement Learning
The multi-task setting can be modeled by casting the environment as a contextual Markov decision process (MDP) ${\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{C},P,R,\gamma,\mu_{0})}$ where $\mathcal{S}\in\mathbb{R}^{N_{\mathcal{S}}}$ and $\mathcal{A}\in\mathbb{R}^{N_{\mathcal{A}}}$ are possibly continuous state and action spaces. $\mathcal{C}$ is a (potentially infinite) set of tasks, with each task represented by an $N_{\mathcal{C}}$-dimensional vector $c\in\mathbb{R}^{N_{\mathcal{C}}}$. $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ models the transition probabilities ($\Delta(\mathcal{S})$ represents the set of probability distributions over $\mathcal{S}$), $R:\mathcal{S}\times\mathcal{C}\to\mathbb{R}$ is a scalar reward function, $\gamma\in(0,1)$ is a discount factor and $\mu_{0}\in\Delta(\mathcal{S})$ is the initial state distribution. In this setting, a policy is simply a state-and-task-conditional action distribution $\bm{\pi}:\mathcal{S}\times\mathcal{C}\to\Delta(\mathcal{A})$ [^1] We use $\bm{\pi}$ and $\pi$ to denote stochastic and deterministic policies, respectively, and $\pi(s,c)$ for realizations.. Any given policy induces a task-conditional distribution over trajectories:
$\bm{\tau}_{\bm{\pi}}\Big{(}\big{(}s_{0},a_{0},s_{1},a_{1},\dots\big{)}\mid c\Big{)}=\mu_{0}(s_{0})\prod_{t=0}^{\infty}\bm{\pi}(a_{t}\mid s_{t},c)\cdot P(s_{t+1}\mid s_{t},a_{t}).$
The discounted returns for a specific task $c\in\mathcal{C}$ or a task distribution $\mu_{c}\in\Delta(\mathcal{C})$ are, respectively,
$J^{\bm{\pi}}_{c}=\mathop{\mathbb{E}}_{(s_{0},\dots)\sim\bm{\tau}_{\bm{\pi}}(c)}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},c)\quad\text{and}\quad J^{\bm{\pi}}_{\mu_{c}}=\mathop{\mathbb{E}}_{c\sim\mu_{c}}J^{\pi}_{c}.$
Reinforcement learning algorithms traditionally aim directly at maximizing $J_{\mu_{c}}^{\bm{\pi}}$, which is notoriously challenging. In the scope of this work, we instead consider an imitation learning setting, in which expert demonstrations from an optimal policy $\pi^{\star}$ are provided. In particular, we focus on behavioral cloning algorithms, which reduce the control problem to a supervised learning problem. Given a set of $N$ task-conditioned, $H$-length trajectories $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ with task labels $c_{1:N}$, behavioral cloning proposes a proxy objective for the policy $\pi$: an empirical estimate of the log-likelihood under the data distribution: $J^{\bm{\pi}}_{\text{proxy}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\log\bm{\pi}(a_{t}^{i}\mid s_{t}^{i},c_{i}).$ If trajectories $\tau_{1:N}$ are obtained from the optimal policy, cover the support of the desired task distribution $\mu_{c}$, and the searched policy class is sufficiently rich, the maximizer of $J^{\bm{\pi}}_{\text{proxy}}$ will also maximize $J^{\bm{\pi}}_{\mu_{c}}$ as $N$ and $H$ increase. However, in general, there is a clear mismatch between $J^{\bm{\pi}}_{\text{proxy}}$ and $J^{\bm{\pi}}_{\mu_{c}}$ (Xu et al., 2020; Maran et al., 2023). Nonetheless, the optimization of $J^{\bm{\pi}}_{\text{proxy}}$ is a relatively straightforward supervised learning problem, while the full RL problem raises several convergence issues, particularly in the offline setting (Levine et al., 2020). Thus, we use $J^{\bm{\pi}}_{\mu_{c}}$ only for evaluation, and carry out optimization through the proxy objective.
3.2 Active Policy Fine-Tuning
In this work, we consider an active fine-tuning scheme for multi-task policies. The goal is to fine-tune a pre-trained policy to perform well on a desired task distribution $\mu_{c}$ using as few expert demonstrations as possible. The agent is allowed $N$ sequential queries for demonstrations according to the fine-tuning budget. The $n$-th query should consist of a task $c_{n}\in\mathcal{C}$. Once the agent selects a task, feedback is received from the optimal policy $\pi^{\star}:\mathcal{S}\times\mathcal{C}\to\mathcal{A}$ (i.e., an optimal demonstrator). At each round the agent receives an $H$-step demonstration conditioned on the chosen task $c_{n}$. This can be seen as a single measurement from a stochastic process over trajectories $\bm{\tau}:\mathcal{C}\to\Delta((\mathcal{S}\times\mathcal{A})^{H})$.
Each observed trajectory up to round $n$ is stored in a dataset $(c_{1:n},\hat{\tau}_{1:n})$, which can be used to fine-tune the policy, and condition the agent’s query at step $n+1$. The process is repeated for $N$ rounds, with the goal of producing a fine-tuned policy that maximizes the expected returns for the desired task distribution $\mu_{c}$.
Modeling assumptions
We take a Bayesian perspective on active multi-task fine-tuning, by assuming a Bayesian model $\bm{\pi}$ over policies. We assume that demonstrations follow a noisy expert: $\bm{\tilde{\pi}}(s,c)=\pi^{\star}(s,c)+\bm{\epsilon}(s,c)$ where $\bm{\epsilon}(s,c)$ is independent noise. We remark, however, that AMF can also be understood from a non-Bayesian perspective as selecting tasks that most quickly minimize the size of frequentist confidence sets around the optimal policy.
4 Method
The active multi-task fine-tuning problem outlined so far requires active data selection for sample-efficient learning. We thus build on top of principled active learning approaches for non-sequential domains (Hübotter et al., 2024), and propose AMF, which selects queries that maximize the expected information gain about the expert policy over its occupancy:
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tau}(c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}).$ (1)
We show in Section 4.1 that, under certain regularity assumptions, the policy learned by AMF converges to the expert policy and matches its performance. These results constitute a first-of-its-kind performance guarantee for active multi-task fine-tuning. The main novelty of this guarantee is the extension of prior work to sequential domains where the visited trajectory $(s_{0},a_{0},s_{1},\dots)\sim\bm{\tau}(c_{n})$ is unknown when selecting the task $c_{n}$ for a demonstration. In Section 4.2, we discuss the design choices that make AMF amenable to optimization in practical settings.
4.1 Performance Guarantees
We begin by presenting the performance guarantees for AMF. Our proof builds upon rates for uncertainty reduction, then ties these to probabilistic convergence guarantees to $\pi^{\star}$, finally resulting in performance guarantees within the MDP. We summarize the main result here, and include a formal proof in Appendix A.
Informal Assumption 1.
We make the following assumptions:
  1. 1.
    The expert policy $\pi^{\star}$ is deterministic, Lipschitz-smooth, lies in the reproducing kernel Hilbert space $\mathcal{H}_{k}(\mathcal{S}\times\mathcal{C})$ of the kernel $k$ with norm $\|\pi^{\star}\|_{k}<\infty$ and induces a Lipschitz-smooth Q-function.
  2. 2.
    The noise $\bm{\epsilon}(s,c)$ affecting demonstrations is conditionally $\rho$-sub-Gaussian and bounded.
  3. 3.
    The dynamics of the contextual MDP $\mathcal{M}$ are Lipschitz-smooth with bounded support, the initial state distribution $\mu_{0}$ has bounded support, and the reward is Lipschitz-smooth.
Under these assumptions, we prove the following performance guarantee for active multi-task behavioral cloning.
Informal Theorem 2 (Performance guarantees for active multi-task BC).
Let all regularity assumptions hold. If each demonstrated task of length $H$ is selected according to the criterion in Equation 1, then with probability $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
$J^{\pi^{\star}}_{\mu_{c}}-J^{\pi_{n}}_{\mu_{c}}\leq O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
where $\pi_{n}$ is the mean of $\bm{\pi}$ at round $n$ and $\gamma_{(Hn)}$ is the maximum information gain about the expert policy from $Hn$ samples, and is sublinear for a large class of problems. The $O(\cdot)$ notation suppresses all multiplicative terms that do not depend on $n$.
Intuitively, this theorem proves that the imitator will eventually achieve the demonstrator’s performance in smooth, regular MDPs with sublinear $\gamma_{(Hn)}$ (for a formal definition, we refer to Lemma 2 in the Appendix). We can also prove a more general result under weaker assumptions: as long as the policy is regular, the imitator will reach the noisy expert performance in arbitrary, non-smooth MDPs, albeit only in expectation. A full derivation of this further result can be found in Appendix B.
4.2 Practical Algorithms
Theorem 2 guarantees that, under regularity assumptions, the adaptive demonstration sampling scheme leads to convergence of the imitator’s performance to the optimal one. However, this criterion involves state occupancies and a conditional entropy term, which are hard to access or estimate in practice. Thus, here we derive a practical objective to be deployed in general settings. We first rephrase the objective from Equation 1 in its entropy form:
$c_{n}$ $=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$ (2)
where we use the definition of mutual information $\mathcal{I}(\cdot|\bm{\tau}(c^{\prime}))=\mathcal{H}(\cdot)-\mathcal{H}(\cdot|\bm{\tau}(c^{\prime}))$, drop the first entropy term as it does not depend on $c^{\prime}$, and rewrite the second entropy term as an expectation over $\bm{\tau}(c^{\prime})$. As long as the task space $\mathcal{C}$ is finite and its cardinality is tractable, the $\operatorname*{arg\,min}$ operator can be evaluated exhaustively, and the expectation over the task distribution $\mu_{c}$ can be computed exactly. When this is not the case, the $\operatorname*{arg\,min}$ can be optimized through discretization, or with sampling-based optimizers. The expectation over $\mu_{c}$ is also not particularly problematic, as it can be computed in closed form (if $\mathcal{C}$ is discrete) or estimated empirically through sampling, as $\mu_{c}$ is assumed to be known. However, two issues need to be resolved: (i) computing the expectation over the noisy expert’s trajectory distribution $\bm{\tau}$, and (ii) estimating the conditional entropy term $\mathcal{H}(\cdot\mid\cdot)$.
Occupancy estimation
Computing the expectation over a policy’s occupancy over states or trajectories is in general intractable in continuous state spaces. Fortunately, a coarse empirical estimate can be obtained as soon as few expert demonstrations become available. The expectation $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}(\cdot)$ can be estimated through a single sample, which is always available in the form of the trajectories $\hat{\tau}_{1:n-1}$ collected so far, as they have effectively been sampled from $\bm{\tau}(c_{1:n-1})$. However, the remaining two expectations (i.e., $\mathop{\mathbb{E}}_{\tau^{\prime}\sim\bm{\tau}(c^{\prime})}(\cdot)$ and $\mathop{\mathbb{E}}_{(s_{0},\dots)\sim\bm{\tau}(c)}(\cdot)$) involve the distribution over trajectories for an arbitrary task, which might not have been demonstrated yet. However, we observe that, at round $n$, the tasks demonstrated so far induce the empirical distribution $\hat{\mu}_{c}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{c_{i}}(\cdot)$, while the trajectories collected similarly induce $\bm{\hat{\tau}}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{\hat{\tau}_{i}}(\cdot)$, where $\delta$ indicates the Dirac delta distribution. We can show that expectations over the trajectory distribution for an arbitrary task $c\in\mathcal{C}$ can be estimated through importance sampling (i.e., by sampling trajectories from $\bm{\hat{\tau}}(\cdot)$ instead of $\bm{\tau}(\cdot|c)$): $\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(\cdot|c)}f(\tau)=\mathop{\mathbb{E}}_{\tau\sim\bm{\hat{\tau}}(\cdot)}\frac{\bm{\tau}(\tau|c)}{\bm{\hat{\tau}}(\tau)}f(\tau)$. The importance weights can then be estimated as
$\frac{\bm{\tau}(\tau|c)}{\bm{\hat{\tau}}(\tau)}$ $\approx\frac{\bm{\tau}(\tau|c)}{\int_{c^{\prime}\in\mathcal{C}}\hat{\mu}_{c}(c^{\prime})\bm{\tau}(\tau|c^{\prime})}=\frac{\bm{\tau}(\tau|c)}{\frac{1}{n-1}\sum_{i=1}^{n-1}\bm{\tau}(\tau|c_{i})}$
$=\frac{(n-1)\mu_{0}(s_{0})\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c)P(s_{t+1}|s_{t},a_{t})}{\sum_{i=1}^{n-1}\mu_{0}(s_{0})\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c_{i})P(s_{t+1}|s_{t},a_{t})}$
$=\frac{(n-1)\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c)}{\sum_{i=1}^{n-1}\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c_{i})}:=w(\tau,c)$ (3)
where $\tau=(s_{0},a_{0},\dots)$ and $\bm{\tilde{\pi}}$ can be approximated with the current estimate of $\bm{\pi}$. Intuitively, the likelihood ratio of a trajectory under two different tasks only depends on the likelihood of actions under the policy, and thus does not require knowledge of the MDP. As the estimate may be inaccurate for small numbers of samples, in practice the algorithm can invest the first few rounds to query a single demonstration for each of the tasks (in case they are countable and few) or to sample the task space uniformly. On the other hand, the high-variance of the estimate can be controlled by practical solutions such as clipping. We present the resulting empirical estimate for Equation 2 in full in Appendix C.
Entropy estimation
The estimation of conditional entropy terms such as $\mathcal{H}(\cdot\mid\cdot)$ has been widely researched in the literature. When the policy is represented through a Gaussian process $GP(\mu,k)$ (Williams & Rasmussen, 2006) with known mean function $\mu$ and kernel $k$, [^2] For simplicity, we consider a single-output GP, but generalize to multi-dimensional policies with multi-output GPs in both experiments and formal proofs. the entropy can be directly quantified by the predicted variance. Let us denote a state-task tuple as $x=(s,c)$, and let $X$ be the sample vector obtained from concatenating states and tasks from previous trajectories (e.g., $c_{1:n-1},\tau_{1:n-1},c^{\prime},\tau^{\prime}$). The unconditional entropy can be measured in closed form as $\mathcal{H}(\bm{\pi}(x))=\frac{1}{2}\log(2\pi k(x,x))+\frac{1}{2}$, and the conditional entropy can be obtained by simply replacing the kernel $k$ with $\hat{k}_{X}(x,x)=k(x,x)-k(x,X)[k(X,X)+\sigma_{\epsilon}^{2}I]^{-1}k(X,x),$ where $\sigma_{\epsilon}^{2}$ is the variance of the observation noise $\bm{\epsilon}(s,c)$, assuming it is distributed according to a zero-mean Gaussian.
Algorithm 1 Practical AMF
Input: initial policy $\pi_{0}$, budget $N$, desired task distr. $\mu_{c}$
Output: fine-tuned policy $\pi_{N}$
Initialize dataset $\mathcal{D}_{0}=\emptyset$
for $n\in[0,\dots,N-1]$ do
     Compute $c_{n}$ as the solution to Equation 2
     Collect new demonstration $\tau_{n}$ for task $c_{n}$
     Update $\pi_{n+1}$ from $\pi_{n}$ with $\mathcal{D}_{n}=\mathcal{D}_{n-1}\cup\{c_{n},\tau_{n}\}$
end for
Thus, when the policy can be modeled as a GP, the only approximation needed concerns occupancy estimation. We refer to this first, practical instantiation as AMF-GP, and present a general algorithmic framework in Algorithm 1. Application of the method to policies parameterized by neural networks will adopt the same scheme. It will also require additional care on three distinct topics: (i) kernel approximations, (ii) batch selection and (iii) forgetting mitigation.
Kernel approximations
When the policy is parameterized through a neural network, estimation of the conditional entropy is far less straightforward. First, we cannot assume the availability of ad-hoc techniques for uncertainty estimation (e.g., Dropout (Srivastava et al., 2014; Gal & Ghahramani, 2016) or ensembles (Lakshminarayanan et al., 2017)), as they might not be featured in pre-trained models. Even if the pre-trained model was perturbed and ensembled for fine-tuning, the ensemble disagreement would not capture the pre-training data distribution. Second, access to pre-training data is in general unrealistic, or hard to manage due to size and ownership of large robotic datasets.
Nevertheless, we can leverage the approximation of neural networks as a linear functions over an embedding space $\bm{\pi}(s,c;\theta)=\bm{\beta}^{\top}\phi_{\theta}(s,c)$, where both weights $\bm{\beta}$ and embeddings $\phi_{\theta}(\cdot)$ exist in a $p$-dimensional latent space (Lee et al., 2019; Khan et al., 2019). This technique does not violate any of the practical constraints listed above, and allows us to adapt the machinery introduced in GP settings. While several embedding strategies exist Jacot et al. (2018); Devlin et al. (2019); Holzmüller et al. (2023), we adopt loss gradient embeddings (Ash et al., 2020a). Assuming the prior $\bm{\beta}\sim\mathcal{N}(0,I)$, the policy $\bm{\pi}(s,c;\theta)$ can be modeled by a Gaussian Process with kernel $k_{\theta}((s,c),(s^{\prime},c^{\prime}))=\langle\phi_{\theta}(s,c),\phi_{\theta}(s^{\prime},c^{\prime})\rangle$. When coupled with this approximation, the conditional entropy objective in Equation 2 can be reformulated:
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}k_{\theta}((s_{t},c),X)[k_{\theta}(X,X)+\sigma_{\epsilon}^{2}I]^{-1}k_{\theta}(X,(s_{t},c)),$ (4)
where $X$ is vector of states and tasks in $(c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$. As the collected dataset grows, the conditioning on previous trajectories $\tau_{1:n-1}$ can instead be addressed by fine-tuning the network’s parameters $\theta$ (e.g., through conventional gradient descent), resulting in updates in the embedding function $\phi_{\theta}$.
Batch selection
Standard practice in deep active learning prescribes the collection of batches of samples at each round (Gal et al., 2017; Sener & Savarese, 2018; Ducoffe & Precioso, 2018). This necessity is partially addressed by the fact that, in our setting, a single demonstration contains several samples for training the policy (as many as the length $H$ of the demonstration). Nevertheless, we can further leverage the GP approximation to select multiple demonstrations at each round. In fact, a simple recursive greedy selection can provide a constant factor approximation on the information gain objective (Krause & Golovin, 2014; Hübotter et al., 2024). In practice, when selecting $M$ demonstrations at the $n$-th round, we can select the $m$-th demonstration out of $M$ by applying the criterion in Equation 1, while making sure to condition entropy estimates on both the $m-1$ demonstrations already selected in this round, and the $M\cdot(n-1)$ demonstrations previously collected.
Dealing with forgetting
A fundamental issue with neural function approximation under shifts to the training distribution is known as forgetting (McCloskey & Cohen, 1989; French, 1999), as any further optimization may catastrophically perturb pre-trained parameters. Common strategies for its mitigation often involve rehearsal (Atkinson et al., 2021; Verwimp et al., 2021) or regularization (Kirkpatrick et al., 2017). Unfortunately, the former is not feasible in this setting due to lack of access to pre-training data, and the latter was not found to be empirically effective (see Appendix G). Scale and a diverse pre-training dataset can also mitigate forgetting (Dyer et al., 2022), but neither can be controlled during fine-tuning. Thus, we opt for a parameter isolating solution (Rusu et al., 2016; Yu et al., 2020), in which the task-space is partitioned (e.g. uniformly), and a copy of the fine-tuning parameters is stored and trained for each partition, thus avoiding negative interference. Inference can then be performed by selecting the parameter set through a nearest-neighbor lookup in task-space. This solution can be easily applied for limited task sets, and can be scaled to large task spaces through discretization schemes. While parameter isolation prevents constructive interference across tasks, we found it to bring a net benefit during fine-tuning.
By combining the approximations required by AMF-GP with three additional design choices (namely kernel approximation, batch selection and parameter isolation), we obtain a method for active multi-task fine-tuning of policies parameterized via neural networks. We refer to this practical instantiation as AMF-NN.
Figure 2: 2D integrator. Starting from the origin, each task involves reaching a given point on a circle, as shown in differently colored trajectories.
5 Experiments
The experiment section is designed to evaluate active multi-task fine-tuning and provide an empirical answer to several questions. We thus reserve a section to each of them.
5.1 When is AMF beneficial?
Uniform pre-training (12/12 tasks)
  AMF-GP    Uniform
Skewed pre-training (6/12 tasks)
Figure 3: Experiments in GP settings for a 2D integrator (see Figure 2). AMF-GP selects tasks that minimize the policy’s posterior entropy and improves the agent’s returns faster than uniform task sampling. In the middle, the improvement in final return over the baseline is greater when the pre-training distribution is skewed and includes fewer tasks. We report return and entropy curves for non-skewed and skewed pre-training (left and right, respectively). We report means and 90% simple bootstrap confidence intervals over 10 random seeds ; dots and crosses mark corresponding measurements.
When none of the assumptions listed in Section 4.1 is violated, AMF is guaranteed to converge to the optimal policy. We furthermore investigate whether AMF also results in faster empirically faster convergence with respect to naive approaches to data collection. To do so, we compare AMF to uniform i.i.d. sampling from the set of tasks $\mathcal{C}$. First, we consider a classic 2D integrator as a benchmark environment (see Figure 2). The agent is a pointmass initialized in the origin, and can directly control its 2D velocity, which is integrated over the past trajectory to return the current state. We can define an infinite task space, in which each task consists of reaching a point on a circle centered on the origin, and the agent is rewarded with the negative Euclidean distance to it. The evaluation distribution $\mu_{c}$ assigns equal probability to 12 points in different directions. The initial state distribution is deterministic, dynamics are both deterministic and smooth, while the expert policy is smooth and corrupted with i.i.d. Gaussian noise. We model the policy as a Gaussian Process with a RBF kernel, and we condition it on a pre-training dataset of $12$ noisy demonstrations. We then collect $50$ additional demonstrations by running both AMF-GP and uniform sampling.
We first consider a perfectly uniform pre-training regime, in which each evaluation task is demonstrated exactly once (Figure 3, left). As the policy’s entropy is minimized, AMF-GP increases the policy’s returns at a higher rate compared to uniform sampling of demonstrations. We then extend this evaluation to several pre-training distributions (Figure 3, middle), and compare the final performance of the two methods as the pre-training budget is allocated to a decreasing number of tasks. As the pre-training distribution becomes more skewed (e.g., when only 6/12 tasks are demonstrated in Figure 3, left), we observe that the performance gap between uniform task sampling and AMF-GP grows larger. This is to be expected, as in this case the information gain from the next demonstration heavily depends on the queried task, and taking the $\operatorname*{arg\,max}$ of the criterion in Equation 1 is significantly better than choosing a random task. Intuitively, in this case, uniform sampling of tasks fails to reliably provide demonstrations for tasks that were observed less often during pre-training.
State-based
0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
Visual
0510Rounds0.200.250.300.350.40Multi-task success rateKitchen, uniform
0510Rounds0.150.200.250.300.350.40Multi-task success rateKitchen, skewed
0510Rounds0.50.60.70.80.9Multi-task success rateMetaworld, uniform
0510Rounds0.40.60.8Multi-task success rateMetaworld, skewed
  AMF-NN ($\sigma^{2}=10^{-4}$)   AMF-NN ($\sigma^{2}=10^{-3}$)    AMF-NN ($\sigma^{2}=10^{-2}$)
  Uniform    Uniform, no parameter isolation
Figure 4: AMF with neural policies in Frankakitchen (top) and Metaworld (bottom). Experiments are repeated for state and RGB inputs (left and right). We evaluate both uniform and skewed pre-training distribution. AMF-NN is overall desirable, and highly beneficial for skewed pre-training distributions. We report means and 90% simple bootstrap confidence intervals over 10 seeds.
5.2 Can AMF scale to high-dimensional tasks?
In realistic settings, the assumptions enabling a formal analysis of AMF are soon violated. As the complexity of the environments of interest increases, most modern behavior cloning applications rely on neural networks for policy parameterization (Reed et al., 2022; Chi et al., 2023). Motivated by this pattern, we now study a second version of our method, AMF-NN, and evaluate its ability to scale to complex, high-dimensional tasks. We consider two common benchmarks for multi-task learning, both with a finite set of tasks.
  • In Metaworld (Yu et al., 2020) we create a scene with a robotic arm, a cup and a faucet, defining 4 tasks: moving the cup to two distinct positions, opening and closing the faucet.
  • In FrankaKitchen (Fu et al., 2020), we consider 5 tasks, namely turning a knob on or off, opening a pivoting or a sliding cabinet, or opening the microwave door.
In both environments, we evaluate AMF-NN when learning from state measurements, as well as from raw pixels. In the first case, the policy is simply parameterized through a MLP, while in the second the MLP receives the embedding of a pre-trained visual encoder (Nair et al., 2022b). The policy is pre-trained on $\approx 15$ total demonstrations, which we allocate either uniformly across all tasks, or only on half of them, reproducing the uniform and skewed regimes from the previous experiments. Afterwards, we apply AMF-NN for $10$ iterations, collecting $2$ demonstrations at each iteration.
Figure 4 reports average multi-task success rates at each iteration compared to a random uniform task selection scheme. For the baseline, we additionally report performance without parameter isolation, highlighting performance degradation due to negative gradient interference across tasks. We observe that AMF’s performance is on par or better than a uniform task sampling scheme when the prior was trained uniformly on all tasks. However, as reported in the previous section, AMF is very beneficial when the pre-training dataset does not uniformly cover the evaluation tasks. These trends are consistent across both environments, and both modalities.
Finally, Figure 4 also highlights the effect of the only hyperparameter introduced by AMF-NN, namely the noise parameter $\sigma^{2}$ in the GP approximation. While selecting an excessively high noise level results in slightly greedier behavior and premature convergence, we overall observe that the method is not particularly sensitive to this hyperparameter, performing fairly reliably across 2 orders of magnitude.
5.3 Is batch-wise task selection important for AMF?
In the GP setting, policy fine-tuning and conditioning for batch-wise selection correspond. Thus, it is not necessary to collect batches of demonstrations. This is not the case for AMF-NN. On one hand, collecting batches of training data is computationally beneficial since training can be parallelized effectively (Sener & Savarese, 2018). On the other hand, each query already returns several samples in our setting, which could make batch selection unnecessary. We thus set out to empirically validate which batch size is desirable for active data collection in our setting.
4.04.55.05.56.06.5Area under success rate curveKitchen, uniform
4.04.55.05.56.06.5Area under success rate curveKitchen, skewed
$B\!=1$ $B\!=2$ $B\!=4$ Uniform
Uniform, no parameter isolation
Figure 5: Influence of batch size $B$ over area under success rate curve with 20 demonstrations.
0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
  Last-layer gradient
  Last-layer    Dropout
Figure 6: AMF performance with alternative uncertainty estimation schemes.
While keeping the total budget fixed to 20 demonstrations, we evaluate AMF-NN with batch sizes $B$ spanning from 1 to 4, and report the area under the success rate curve in Figure 5. For convenience, we report the same metric for the uniform selection baselines from Figure 4, for which the data selection strategy does not depend on the batch size. We find that larger batch sizes are not necessarily desirable.
5.4 How do uncertainty estimates for AMF compare?
As entropy estimation is at the core of AMF-NN, we additionally compare the adopted GP approximation with loss-gradient embeddings to other approaches from the literature. In praticular, we also consider an alternative GP approximation using last-layer embeddings (Holzmüller et al., 2023), as well as test-time Dropout (Loquercio et al., 2020). For the latter, each batch is simply filled with demonstrations from the task maximizing prior entropy, that is $\operatorname*{arg\,max}_{c\in\mathcal{C}}\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;(s_{0},\dots)\sim\bm{\tau}(c)}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid\tau_{1:n-1})$. Both of these schemes are in practice desirable, as they do not require access to action labels. However, we observe that these two schemes are prone to early convergence to suboptimal task choices, or are less effective in driving task selection. Hence, as shown in Figure 6, multi-task performance is in general lower, suggesting that the entropy estimation technique is crucial to AMF-NN.
5.5 Can AMF be applied to off-the-shelf models?
  AMF-NN    Uniform
  Uniform, no parameter isolation
024Rounds0.400.450.500.550.600.65Multi-task success rateWidowX
Figure 7: Evaluation on life-like WidowX tasks. AMF-NN can be applied to large-scale settings.
As AMF-NN has minimal requirements (essentially, access to a differentiable pre-trained prior is sufficient), it should be widely applicable. In this section, we investigate scaling our evaluation to recently published open-source generalist policies. For this purpose, we chose Octo (Octo Model Team et al., 2024). This model relies on a transformer backbone for integrating multimodal information (in the form of state sensors, camera images and text or RGB task descriptions), and uses a diffusion-based policy head for action prediction (Chi et al., 2023). For computational reasons, we will focus on fine-tuning the action head alone. Octo is pre-trained on a large-scale real-world robotic dataset (Collaboration, 2023), and is thus designed for inference on physical hardware. Nonetheless, a recently proposed evaluation suite enables simulated evaluations that statistically correlate with real-world results (Li et al., 2024). We thus collect rollouts from a pre-trained Octo on the WidowX tasks, and filter them to only include successes, akin to self-distillation schemes Bousmalis et al. (2024). On availability of such self-supervised demonstrations, we then apply AMF-NN for $4$ iterations, providing $4$ demonstrations in each round. The results are reported in Figure 7.
As all evaluation tasks are largely demonstrated in the pre-training dataset (Collaboration, 2023), we find that AMF-NN does not improve significantly upon uniform task collection, confirming the trend we observed for uniform pre-training distributions in Figure 4. Nonetheless, we observe that it constitutes an effective method for data selection, and can be applied as a drop-in replacement for fine-tuning of off-the-shelf models. [^3] This evaluation also reports an interesting trend, that is a vast reduction in catastrophic forgetting, to the point that parameter isolation is not necessary. This anecdotal evidence can be seen as an instance of a general trend of mitigated catastrophic forgetting in large models (Dyer et al., 2022).
6 Discussion
As generalist robotic policies gain prominence, a new set of challenges and opportunities emerge. This work responds to this trend by investigating an active multi-task fine-tuning scheme, which adaptively selects the task to be demonstrated for sample-efficient multi-task behavioral cloning. This approach is developed from first principles, extending a formally-motivated, information-based criterion to trajectories over dynamic systems. The resulting method is both formally supported by novel performance guarantees and widely applicable. Moreover, a practical instantiation enables sample-efficient multi-task fine-tuning across GP and neural network policy classes.
Naturally, active multi-task fine-tuning has several limitations. When coupled with neural networks, the algorithm relies on uncertainty estimation techniques, which remain an open problem. While the approximation we leverage is informative in our experiments, AMF could benefit if large pre-trained policies would allow other off-the-shelf uncertainty quantification techniques (e.g., through model ensembling during pre-training). Second, we found the performance of AMF to depend naturally on the pre-training data distribution. While AMF induces efficient learning for skewed pre-training distributions, it naturally brings more modest gains when the pre-trained policy is equally capable for all tasks, and uniform task sampling is sufficient.
On top of addressing the current limitations, this work suggests multiple interesting directions. An extensive empirical evaluation of active fine-tuning with large-scale generalist policies is clearly desirable, but remains infeasible at the moment due to the scarce availability of open-source models and benchmarks. Another future research direction would involve direct estimation of the RL objective, thus removing the dependence on non-equivalent BC proxy objectives.
Acknowledgments
We thank Nico Gürtler, Ji Shi, Andreas René Geist and Usman Namjad for the valuable discussions. Marco Bagatella is supported by the Max Planck ETH Center for Learning Systems. Georg Martius is a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. We acknowledge the support from the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039B), from the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943, and from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545.
Reproducibility Statement
Appendix I describes implementation details and hyperparameters across all experiments. We additionally open-source a clean and commented implementation of AMF-GP and AMF-NN at the anonymous website.
Appendix A Performance guarantees under regularity assumptions
This sections retrieves guarantees on the performance of the imitator policy as a function of the number of provided demonstrations $n$. At first, this analysis focuses on policies over a single-dimensional action space. An extension to multi-dimensional outputs is introduced later on. The general sketch of the proof can be informally described as follows:
  • we first introduce the regularity assumptions required for the guarantees;
  • we then show that, in Lipschitz, bounded MDPs, the effect of stochasticity on information gain at each round can be controlled;
  • we show how the variance over the imitator’s policy shrinks according to the maximum information gain at each round, which in turn depends on the maximum information gain over a set of queries to the expert;
  • starting from the previous result, we leverage a well-known theorem (Abbasi-Yadkori, 2013) to retrieve a probabilistic, anytime guarantee on the error of the imitator;
  • we quantify the relationship between the imtator’s error and its performance, thus retrieving our main theoretical result.
A.1 Assumptions
It is clear that bounding imitation performance would be hopeless without any regularity assumption, as slight errors in the imitator’s policy could result in arbitrary differences in return. We thus introduce the following:
Assumption 1.
(Regular, noisy policy) We assume that the optimal policy $\pi^{\star}\sim GP(\mu,k)$ with known mean function $\mu$ and kernel $k$. Furthermore the noise $\bm{\epsilon}(s,c)$ is mutually independent and zero-mean Gaussian, with known variance $\rho^{2}(s,c)>0$ for all $(s,c)\in\mathcal{S\times C}$.
In order to motivate further assumptions, let us recall the criterion from Equation 1:
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tau}(c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}).$ (5)
Through this section, we will use a slightly more precise formulation:
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\tau\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}),$ (6)
in which we clarify that the mutual information is only computed with respect to the actions of the noisy expert, and overload the notation with $\bm{\tilde{\pi}}(\tau,c^{\prime})=(\bm{\tilde{\pi}}(s_{i},c^{\prime}))_{0}^{H-1}$ for $\tau=(s_{i},a_{i})_{0}^{H-1}$. The criterion selects the task $c_{n}$ with the greatest expected mutual information between the policy and the trajectory associated with the task. We note that, the objective produces a fully deterministic sequence of tasks, as all stochasticity is resolved in the expectation. Nevertheless, the actual sequence of states at which the demonstrator is queried remains stochastic. For this reason, we require the following two sets of assumptions to ensure that information gained along empirical trajectories is not arbitrarily smaller than the expected one.
Assumption 2.
(Lipschitz, bounded MDP and policy) Given the contextual MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{C},P,R,\gamma,\mu_{0})$ and the noisy expert $\bm{\tilde{\pi}}$ we assume that, for every $\{(s,c,a),(s^{\prime},c^{\prime},a^{\prime})\}\subseteq\mathcal{S\times C\times A}$:
  • the support of the initial state distribution $\mu_{0}$ is bounded by an $\epsilon_{\mu_{0}}$-ball
    $\max_{s_{l},s_{h}\in\text{supp}(\mu_{0})}\|s_{h}-s_{l}\|_{2}\leq\epsilon_{\mu_{0}},$
  • the transition kernel $P$ is $L_{P}$-smooth
    $\mathcal{W}(P(\cdot|s,a),P(\cdot|s^{\prime},a^{\prime}))\leq L_{P}\cdot d((s,a),(s^{\prime},a^{\prime})),$
    where $d((s,a),(s^{\prime},a^{\prime}))=\|s-s\|_{2}+\|a-a^{\prime}\|_{2}$, and $\mathcal{W}(\cdot,\cdot)$ is the Wasserstein $1$-distance with respect to $d(\cdot,\cdot)$; furthermore, the support of $P(\cdot|s,a)$ is bounded by an $\epsilon_{P}$-ball
    $\max_{s_{l},s_{h}\in\text{supp}(P(\cdot|s,a))}\|s_{h}-s_{l}\|_{2}\leq\epsilon_{P},$
  • the reward function $R$ is $L_{R}$-smooth
    $|R(s,c,a)-R(s^{\prime},c^{\prime},a^{\prime})|<L_{R}\cdot d((s,c,a),(s^{\prime},c^{\prime},a^{\prime})),$
    where $d((s,c,a),(s^{\prime},c^{\prime},a^{\prime}))=\|s-s\|_{2}+\|c-c^{\prime}\|_{2}+\|a-a^{\prime}\|_{2}$,
  • the noisy expert $\bm{\tilde{\pi}}$ is $L_{\pi}$-smooth
    $\mathcal{W}(\bm{\tilde{\pi}}(\cdot|s,c,a),P(\cdot|s^{\prime},c^{\prime},a^{\prime}))\leq L_{\pi}\cdot d((s,c,a),(s^{\prime},c^{\prime},a^{\prime})),$
    where $d((s,c,a),(s^{\prime},c^{\prime},a^{\prime}))=\|s-s\|_{2}+\|c-c^{\prime}\|_{2}+\|a-a^{\prime}\|_{2}$ and $\mathcal{W}(\cdot,\cdot)$ is the $1$-Wasserstein distance with respect to $d(\cdot,\cdot)$; furthermore, the support of $\bm{\tilde{\pi}}(\cdot|s,a)$ is bounded by an $\epsilon_{\pi}$-ball
    $\max_{a_{l},a_{h}\in\text{supp}(\bm{\tilde{\pi}}(\cdot|s,c))}\|a_{h}-a_{l}\|_{2}\leq\epsilon_{P},$
  • finally, the Q-function for the expert $\pi^{\star}$ is $L_{Q}$-smooth
    $|Q^{\pi^{\star}}(s,c,a)-Q^{\pi^{\star}}(s^{\prime},c^{\prime},a^{\prime})|\leq L_{Q}\cdot d((s,c,a),(s^{\prime},c^{\prime},a^{\prime})).$
We note that smoothness of the noisy expert is guaranteed by construction if the expert $\pi^{\star}$ is $L_{\pi}$-smooth.
Assumption 3.
(Smooth MI) For every pair of sequences of trajectories $\{\tau_{1:n-1},\tau^{\prime}_{1:n-1}\}\subseteq(\mathcal{S\times A})^{H(n-1)}$, $(s,c)\in\mathcal{S\times C}$, $c_{1:n-1}\in\mathcal{C}^{n-1}$, $\tilde{\tau}\in(\mathcal{S\times A})^{H}$ and $c_{n}\in\mathcal{C}$, we assume that the mutual information at step $n$ is $L_{I}$-smooth with respect to the mean square deviation of collected trajectories:
$|\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tilde{\tau},c_{n})|c_{1:n-1},\tau_{1:n-1})-\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tilde{\tau},c_{n})|c_{1:n-1},\tau^{\prime}_{1:n-1})|\leq L_{I}\cdot d(\tau_{1:n-1},\tau^{\prime}_{1:n-1}),$
where $d((s_{0,1},a_{0,1},\dots s_{H,n-1},a_{H,n-1}),(s^{\prime}_{0,1},a^{\prime}_{0,1},\dots s^{\prime}_{H,n-1},a^{\prime}_{H,n-1})=\frac{1}{n-1}\sum_{m=1}^{n-1}(\sum_{t=0}^{H-1}$ $\|s_{t,m}-s^{\prime}_{t,m}\|_{2}^{2}+\|a_{t,m}-a^{\prime}_{t,m}\|_{2}^{2})^{\frac{1}{2}}$ is the mean square deviation over the concatenation of trajectories.
A.2 Proof
We first prove that, under Assumptions 2 and 3, the effect of stochasticity on the mutual information at step $n$ is bounded.
Lemma 1.
Let Assumptions 2 and 3 hold. Fix a sequence of tasks $c_{1:n-1}$ and consider two arbitrary sequences of trajectories $\tau_{1:n-1}$ and $\tau^{\prime}_{1:n-1}$ sampled from $\bm{\tau}(c_{1:n-1})$. Fix one state-task pair $(s,c)\in\mathcal{S\times C}$, one task $c_{n}\in\mathcal{C}$ and one trajectory $\tilde{\tau}\sim\bm{\tau}(c_{n})$. Let $\epsilon_{n}=8H^{\frac{3}{2}}(1+\max(L_{P},L_{\pi}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$. The difference in mutual information when conditioning on the two sequences of trajectories can be bounded:
$|\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tilde{\tau},c_{n})|\tau_{i:n-1})-\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tilde{\tau},c_{n})|\tau^{\prime}_{i:n-1})|\leq\epsilon_{n}.$
Proof.
Under Assumption 3 it is sufficient to show that stochasticity in the MDP does not cause the demonstrator’s trajectories to deviate excessively. This is a direct consequence of smoothness and boundedness, which we assume in Assumption 2, and can be shown by induction. Let us fix a task $c_{n}\in\mathcal{C}$ and consider two trajectories $\tau,\tau^{\prime}\sim\bm{\tau}(c_{n})$. For the two initial states $(s_{0},s^{\prime}_{0})$, boundedness of the initial state distribution $\mu_{0}$ implies that $\|s_{0}-s^{\prime}_{0}\|_{2}\leq\epsilon_{\mu_{0}}$. Now, assuming that the distance between two states $(s_{t},s^{\prime}_{t})$ is bounded as $\|s_{t}-s^{\prime}_{t}\|_{2}\leq\epsilon_{t}$, we have that
$\epsilon_{t+1}$ $:=\|s_{t+1}-s^{\prime}_{t+1}\|_{2}$ (7)
$\stackrel{{\scriptstyle(i)}}{{\leq}}\mathcal{W}(P(\cdot|s_{t},a_{t}),P(\cdot|s^{\prime}_{t},a^{\prime}_{t}))+2\epsilon_{P}$ (8)
$\leq L_{P}\cdot(\|s_{t}-s^{\prime}_{t}\|_{2}+\|a_{t}-a^{\prime}_{t}\|_{2})+2\epsilon_{P}$ (9)
$=L_{P}\cdot(\epsilon_{t}+\|a_{t}-a^{\prime}_{t}\|_{2})+2\epsilon_{P}$ (10)
$\stackrel{{\scriptstyle(ii)}}{{\leq}}L_{P}\cdot(\epsilon_{t}+\mathcal{W}(\bm{\tilde{\pi}}(\cdot|s_{t},c_{t}),\bm{\tilde{\pi}}(\cdot|s^{\prime}_{t},c^{\prime}_{t}))+2\epsilon_{\pi})+2\epsilon_{P}$ (11)
$\leq L_{\pi}\cdot(\epsilon_{t}+L_{\pi}\cdot\|s_{t}-s^{\prime}_{t}\|_{2}+2\epsilon_{\pi})+2\epsilon_{P}$ (12)
$=L_{P}\cdot(\epsilon_{t}+L_{\pi}\cdot\epsilon_{t}+2\epsilon_{\pi})+2\epsilon_{P}$ (13)
$=L_{P}\cdot((L_{\pi}+1)\cdot\epsilon_{t}+2\epsilon_{\pi})+2\epsilon_{P}$ (14)
$=L_{P}(1+L_{\pi})\epsilon_{t}+2(L_{P}\epsilon_{\pi}+\epsilon_{P})$ (15)
$:=A\epsilon_{t}+B,$ (16)
where Lemma 9 was used in (i) and (ii); Assumption 2 and the fact that $c_{t}=c^{\prime}_{t}$ were used through the rest of the derivation. The recurrence relation can be easily unrolled as
$\epsilon_{t}$ $\leq A^{t}\epsilon_{0}+\sum_{i=0}^{t-1}A^{i}B$ (17)
$\leq A^{t}\epsilon_{0}+\max(A^{t-1},1)Bt$ (18)
$\leq\max(A,1)^{t}(\epsilon_{0}+Bt)$ (19)
$=\max(L_{P}(1+L_{\pi}),1)^{t}(\epsilon_{0}+2t(L_{P}\epsilon_{\pi}+\epsilon_{P}))$ (20)
$\leq(1+L_{P})^{t}(1+L_{\pi})^{t}(\epsilon_{0}+2t((1+L_{P})\epsilon_{\pi}+\epsilon_{P}))$ (21)
$\leq(1+L_{P})^{t}(1+L_{\pi})^{t}(2t(1+L_{P})\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P}))$ (22)
$=2t(1+L_{P})^{t+1}(1+L_{\pi})^{t}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P}),$ (23)
thus bounding the L2 distances between states at each step of the trajectory $\epsilon_{t}=\|s_{t}-s^{\prime}_{t}\|_{2}$. We note that the distance between actions can also be easily bound by Lemma 9: $\|a_{t}-a^{\prime}_{t}\|_{2}\leq L_{\pi}\epsilon_{t}+2\epsilon_{\pi}$. This can in turn be related to distances over trajectories. Let us fix $c_{1:n-1}\in\mathcal{C}$ and consider $\tau_{1:n-1},\tau^{\prime}_{1:n-1}\sim\bm{\tau}(c_{1:n-1})$. We have that
$d(\tau_{1:n-1},\tau^{\prime}_{1:n-1})$ $=\frac{1}{n-1}\sum_{m=1}^{n-1}(\sum_{t=0}^{H-1}\|s_{t,m}-s^{\prime}_{t,m}\|_{2}^{2}+\|a_{t,m}-a^{\prime}_{t,m}\|_{2}^{2})^{\frac{1}{2}}$ (24)
$\leq(\sum_{t=0}^{H-1}\epsilon^{2}_{t}+(L_{\pi}\epsilon_{t}+2\epsilon_{\pi})^{2})^{\frac{1}{2}}$ (25)
$\leq(\sum_{t=0}^{H-1}\epsilon^{2}_{t}+(L_{\pi}\epsilon_{t}+\epsilon_{t})^{2})^{\frac{1}{2}}$ (26)
$=(\sum_{t=0}^{H-1}\epsilon^{2}_{t}+(1+L_{\pi})^{2}\epsilon_{t}^{2})^{\frac{1}{2}}$ (27)
$\leq(\sum_{t=0}^{H-1}2(1+L_{\pi})^{2}\epsilon_{t}^{2})^{\frac{1}{2}}$ (28)
$=(2(1+L_{\pi})^{2}\sum_{t=0}^{H-1}\epsilon_{t}^{2})^{\frac{1}{2}}$ (29)
$=\sqrt{2}(1+L_{\pi})(\sum_{t=0}^{H-1}\epsilon_{t}^{2})^{\frac{1}{2}}$ (30)
$\leq\sqrt{2}(1+L_{\pi})(H\epsilon_{H-1}^{2})^{\frac{1}{2}}$ (31)
$=\sqrt{2H}(1+L_{\pi})\epsilon_{H-1}$ (32)
$\leq\sqrt{2H}(1+L_{\pi})\cdot 2(H-1)(1+L_{P})^{H}(1+L_{\pi})^{H-1}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$ (33)
$\leq 4H^{\frac{3}{2}}(1+L_{P})^{H}(1+L_{\pi})^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$ (34)
$\leq 8H^{\frac{3}{2}}(1+\max(L_{P},L_{\pi}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P}).$ (35)
Having obtained an upper bound on the distance between sequences of trajectories, the result follows naturally from smoothness of mutual information according to Assumption 3.
We can now focus on the main result. We start by introducing an important measure, quantifying the maximum information gain at each round:
$\Gamma_{n}:=\max_{c^{\prime}\in\mathcal{C}}\psi_{n}(c^{\prime})=\max_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau^{\prime},c^{\prime})\mid c_{1:n-1},\tau_{1:n-1})$ (36)
We note that the criterion in Equation 6 takes the $\operatorname*{arg\,max}$ of the same quantity $\Gamma_{n}$ maximizes over. As common in the literature (Bogunovic et al., 2016; Kothawade et al., 2020; Hübotter et al., 2024), we make a standard assumption on diminishing informativeness.
Assumption 4.
For each $n,i\in\mathbb{N}$ with $i\leq n$, the maximum information gain at round $n$ is not greater than the maximum information gain at round $i$:
$\Gamma_{n}\leq\Gamma_{i}.$
This can be leveraged to show that the expected mutual information is sublinear in the number of rounds $n$. From this point, we overload the notation and allow policies (e.g. $\bm{\pi}$) to map vector to random vectors, that is $\bm{\pi}((x_{0},\dots,x_{n-1}))=(\bm{\pi}(x_{0}),\dots,\bm{\pi}(x_{n-1}))$ for $(x_{0},\dots,x_{n-1})\in(\mathcal{S\times C})^{n}$.
Lemma 2.
Under Assumptions 1 and 4, if $(c_{0},\dots,c_{n})$ follows the criterion in Equation 6, then $\Gamma_{n}\leq\frac{H}{n}\gamma_{(Hn)}$, where $\gamma_{(Hn)}=\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|\leq Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$.
Proof.
$\Gamma_{n}$ $=\frac{1}{n}\sum_{i=0}^{n-1}\Gamma_{n}$ (37)
$\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{n}\sum_{i=0}^{n-1}\Gamma_{i}$ (38)
$=\frac{1}{n}\sum_{i=0}^{n-1}\max_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{0},c);\bm{\tilde{\pi}}(\tau^{\prime},c^{\prime})\mid c_{1:n-1},\tau_{1:n-1})$ (39)
=$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{n}\sum_{i=0}^{n-1}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c_{n})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{0},c);\bm{\tilde{\pi}}(\tau^{\prime},c_{n})\mid c_{1:n-1},\tau_{1:n-1})$ (40)
$=\frac{1}{n}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{n}\sim\bm{\tau}(c_{n})\\ \tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\end{subarray}}\sum_{i=0}^{n-1}\mathcal{I}(\bm{\pi}(s_{0},c);\bm{\tilde{\pi}}(\tau_{n},c_{n})\mid c_{1:n-1},\tau_{1:n-1})$ (41)
=$\stackrel{{\scriptstyle(iii)}}{{=}}\frac{1}{n}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{n}\sim\bm{\tau}(c_{n})\\ \tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\end{subarray}}\mathcal{I}(\bm{\pi}(s_{0},c);\bm{\tilde{\pi}}(\tau_{1:n},c_{1:n}))$ (42)
$\leq\frac{1}{n}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|=Hn\end{subarray}}\mathcal{I}(\bm{\pi}(s_{0},c);\bm{\tilde{\pi}}(X))$ (43)
$\leq\frac{1}{n}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|=Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ (44)
$=\frac{H}{n}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|=Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ (45)
$=\frac{H}{n}\gamma_{(Hn)}$ (46)
where (i) follows from Assumption 4, (ii) follows from Equation 6, (iii) is due to the chain rule of mutual information. We note that $\gamma_{n}=\max_{X\subseteq\mathcal{S}\times\mathcal{C},\;|X|\leq n}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ is sublinear for a large class of GPs. In this cases, a looser upper bound would be $H^{2}\frac{\gamma_{n}}{n}$. ∎
This bound on expected round-wise mutual information can then be leveraged to describe how the total variance shrinks over rounds.
Lemma 3.
(Uniform convergence of marginal variance, following Hübotter et al. (2024)) Under Assumption 1, 2 and 3, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
$\sigma_{n}^{2}(s,c)\leq(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau_{\text{min}}^{2}},$
where $\bar{\sigma}^{2}=\max_{(s,c)\in\mathcal{S\times C}}\sigma^{2}_{0}(s,c)+\rho^{2}(s,c)$ and $\tau_{\text{min}}=\min_{s,c\in\mathcal{S\times C}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}$.
Proof.
$\sigma_{n}^{2}(s,c)$ $=\mathrm{Var}[\bm{\pi}(s,c)\mid c_{1:n},\tau_{1:n}]$ (47)
$=\big{(}\mathrm{Var}[\bm{\pi}(s,c)\mid c_{1:n},\tau_{1:n}]+\rho^{2}(s,c)\big{)}-\rho^{2}(s,c)$ (48)
$=\mathrm{Var}[\bm{\tilde{\pi}}(s,c)\mid c_{1:n},\tau_{1:n}]-\mathrm{Var}[\bm{\tilde{\pi}}(s,c)\mid\bm{\pi}(s,c),c_{1:n},\tau_{1:n}]$ (49)
$\stackrel{{\scriptstyle(i)}}{{\leq}}\bar{\sigma}^{2}\log\left(\frac{\mathrm{Var}[\bm{\tilde{\pi}}(s,c)\mid c_{1:n},\tau_{1:n}]}{\mathrm{Var}[\bm{\tilde{\pi}}(s,c)\mid\bm{\pi}(s,c),c_{1:n},\tau_{1:n}]}\right)$ (50)
$=2\bar{\sigma}^{2}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(s,c)\mid c_{1:n},\tau_{1:n})$ (51)
$=2\bar{\sigma}^{2}\frac{1}{\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(s,c)\mid c_{1:n},\tau_{1:n})$ (52)
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(s,c)\mid c_{1:n},\tau_{1:n})$ (53)
$\leq\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (54)
$\leq\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (55)
$\leq\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}}\frac{1}{\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\mathbf{1}_{s\in(s_{0},\dots)}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\mathbf{1}_{s\in(s_{0},\dots)}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (56)
$\leq\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\mathbf{1}_{s\in(s_{0},\dots)}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (57)
$=\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\mathbf{1}_{s\in(s_{0},\dots)}\mathcal{I}(\bm{\pi}(s,c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (58)
$\leq\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\mathbf{1}_{s\in(s_{0},\dots)}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (59)
$\leq\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (60)
$\stackrel{{\scriptstyle(iii)}}{{\leq}}(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}^{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (61)
$=(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}^{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau\sim\bm{\tau}(c)\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (62)
$\leq(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}}{\tau_{\text{min}}^{2}}\max_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau^{\prime},c^{\prime})\mid c_{1:n},\tau_{1:n})$ (63)
$=(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau_{\text{min}}^{2}}.$ (64)
where (i) follows from Lemma 8 and monotonicity of variance, (ii) holds as the state $s$ is within the support of $\bm{\tau}(\cdot|c)$, and (iii) follows from Lemma 1 as the difference between the expected mutual information and the mutual information for a realized trajectory is less than the difference in mutual information for two arbitrary realized trajectories. ∎
This result can then be translated to the agnostic setting, for a regular policy $\pi^{\star}$, which we still model through the stochastic process $\bm{\pi}$. Without loss of generality we will assume that the prior variance is bounded by $\mathrm{Var}[\bm{\pi}(s,c)]\leq 1$.
Lemma 4.
(Well-calibrated confidence intervals, following Abbasi-Yadkori (2013)) Pick $\delta\in(0,1)$. Assume that $\pi^{\star}$ lies in the RKHS $\mathcal{H}_{k}(\mathcal{C})$ of the kernel $k$ with norm $||\pi^{\star}||_{k}<\infty$, the noise $\epsilon_{n}$ is conditionally $\rho$-sub-Gaussian, and $\gamma_{n}$ is sublinear in $n$. Let $\beta_{n}(\delta)=\|\pi^{\star}\|_{k}+\rho\sqrt{2(\gamma_{(Hn)}+1+\log(1/\delta))}$. Then, for any $n>1$ and $(s,c)\in\mathcal{S\times C}$, $GP(\mu_{n},k)$ is an all-time well-calibrated model of $\pi^{\star}$. Thus, jointly with probability at least $1-\delta$,
$|\pi^{\star}(s,c)-\mu_{n}(s,c)|\leq\beta_{n}(\delta)\sigma_{n}.$
We note that $\beta_{n}(\delta)$ depends on $\gamma_{(Hn)}$ as $Hn$ samples from the demonstrator’s policy are collected up to round $n$. Combining Lemmas 3 and 4 we easily get for all $(s,c)\in\mathcal{S\times C}$ and $n\geq 0$ with probability $1-\delta$:
[…]
While the analysis has so far dealt with a scalar $\pi^{\star}$, a simple union bound can guarantee that
$\|\pi^{\star}(s,c)-\mu_{n}(s,c)\|_{1}\leq\beta^{\prime}_{n}(\delta)\|\bar{\sigma}\|_{1}\Big{(}(1+\epsilon_{n})\frac{2\Gamma_{n}}{\tau_{\text{min}}^{2}}\Big{)}^{\frac{1}{2}}$ (66)
with probability at least $1-\delta$ for an action space of dimension $|\mathcal{A}|$, where now $\beta^{\prime}_{n}(\delta)=\|\pi^{\star}\|_{k}+\rho\sqrt{2(\gamma_{(Hn)}+1+\log(|\mathcal{A}|/\delta))}$. From now on, we will refer to $\mu_{n}$ as $\pi_{n}$. We are thus able to globally bound the $L_{1}$ distance of the imitator policy with respect to the expert policy with high probability under active fine-tuning.
It is clear that, even if this distance is small, the performance of an imitator which does not exactly match the expert ($\pi_{n}(s,c)\neq\pi^{\star}(s,c)$ for some $(s,c)\in\mathcal{S}\times\mathcal{C})$ can be arbitrarily low for arbitrary MDPs. It is however possible to show that, as long as the Q-function of the expert is smooth, the performance gap to the expert can be controlled. We note that, in case $\gamma L_{P}(1+L_{\pi^{\star}})<1$, then the Q-function $Q^{\pi^{\star}}$ is guaranteed to be $L_{Q}$-Lipschitz continuous with $L_{Q}\leq\frac{L_{R}}{1-\gamma L_{P}(1+L_{\pi^{\star}})}$ (Rachelson & Lagoudakis, 2010). If smoothness holds, it is easy to connect the divergences in action space to performance gaps (Maran et al., 2023).
Lemma 5.
Let $\pi$ and $\pi^{\prime}$ denote two deterministic policies. If the state-action value function $Q^{\pi^{\prime}}$ is $L_{Q^{\pi^{\prime}}}$-Lipschitz continuous, then:
$|J^{\pi}-J^{\pi^{\prime}}|\leq\frac{L_{Q^{\pi^{\prime}}}}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}[\|\pi^{\prime}(s,c)-\pi(s,c)\|_{1}].$
Proof.
Given a function $f:\mathcal{A}\to\mathbb{R}$, we denote the Lipschitz semi-norm $\|f(\cdot)\|_{L}=\sup_{a,a^{\prime}\in\mathcal{A}}\frac{|(f(a)-f(a^{\prime})|}{\|a-a^{\prime}\|_{2}}$. We have:
$J^{\pi}-J^{\pi^{\prime}}$ =$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\mathbb{E}_{a\sim\pi(\cdot\mid s,c)}[A^{\pi^{\prime}}(s,c,a)]\Bigg{]}$ (67)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\mathbb{E}_{a\sim\pi(\cdot\mid s)}[Q^{\pi^{\prime}}(s,c,a)]-V^{\pi^{\prime}}(s,c)\Bigg{]}$ (68)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\int_{a\in\mathcal{A}}\pi(a\mid s)Q^{\pi^{\prime}}(s,c,a)-V^{\pi^{\prime}}(s,c)\Bigg{]}$ (69)
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\int_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,c,a)[\pi(a\mid s,c)-\pi^{\prime}(a\mid s,c)]\Bigg{]}$ (70)
$\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\int_{a\in\mathcal{A}}\sup_{s,c\in\mathcal{S\times C}}Q^{\pi}(s,c,a)[\pi(a\mid s,c)-\pi^{\prime}(a\mid s,c)]\Bigg{]}$ (71)
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}\Bigg{[}\|\sup_{s,c\in\mathcal{S\times C}}Q^{\pi^{\prime}}(s,c,\cdot)\|_{L}\mathcal{W}(\pi(\cdot\mid s,c),\pi^{\prime}(\cdot\mid s,c))\Bigg{]}$ (72)
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\frac{L_{Q^{\pi^{\prime}}}}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}}[\mathcal{W}(\pi(\cdot\mid s,c),\pi^{\prime}(\cdot\mid s,c))]$ (73)
=$\stackrel{{\scriptstyle(iv)}}{{=}}\frac{L_{Q^{\pi^{\prime}}}}{1-\gamma}\mathbb{E}_{s,c\sim d^{\pi}}[\|\pi(\cdot\mid s,c)-\pi^{\prime}(\cdot\mid s,c)\|_{1}]$ (74)
where (i) follows from the performance difference lemma (Kakade & Langford, 2002), (ii) follows from the definition of $L_{1}$ Wasserstein distance, (iii) holds as $L_{Q^{\pi}}\geq\|\sup_{s,c\in\mathcal{S\times C}}Q^{\pi}(s,c,\cdot)\|_{L}$ and (iv) follows from both policies being deterministic. The proof is completed by taking the absolute value on both sides. ∎
So far, we have shown rates of convergence for the imitator, and connected its error to performance. Our main formal result can be shown by coordinating the lemmas so far presented.
Theorem 3.
(Performance guarantees for active multi-task BC) Let Assumptions 2, 3 and 4 hold. Pick $\delta\in(0,1)$. Assume that $\pi^{\star}$ lies in the RKHS $\mathcal{H}_{k}(\mathcal{C})$ of the kernel $k$ with norm $||\pi^{\star}||_{k}<\infty$, the noise $\epsilon_{n}$ is conditionally $\rho$-sub-Gaussian, and $\gamma_{n}$ is sublinear in $n$. If each demonstrated task is selected according to the criterion in Equation 1, then with probability at least $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
$J^{\pi^{\star}}-J^{\pi_{n}}\leq\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}=O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
where $\epsilon_{n}=8H^{\frac{3}{2}}(1+\max(L_{\pi},L_{P}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$. Furthermore, if $\gamma_{n}=O(\log n)$ (e.g., for linear kernels), then $J^{\pi^{\star}}-J^{\pi}\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$.
Proof.
$J^{\pi^{\star}}-J^{\pi}$ =$\stackrel{{\scriptstyle(i)}}{{=}}|J^{\pi}-J^{\pi^{\star}}|$ (75)
[…]
$=\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}$ (78)
where (i) is due to the fact that $J^{\pi^{\star}}\geq J^{\pi}$ for any policy $\pi$, and the expectation fades due to uniform convergence. The only terms with a dependency on $n$ are $\beta^{\prime}_{n}(\delta)=O(\gamma_{(Hn)}^{\frac{1}{2}})$ and $\Gamma_{n}=O(\gamma_{(Hn)}n^{-1})$, which can be combined in the asymptotic notation in the Theorem. If $\gamma_{n}=O(\log n)$, then $J^{\pi^{\star}}-J^{\pi}=O(n^{-\frac{1}{2}}\log n)\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$. For a summary of magnitudes of $\gamma_{n}$ for common kernels, we refer to Table 3 in Hübotter et al. (2024). ∎
Appendix B Guarantees in non-Lipschitz MDPs
The main result reported in Theorem 3 provides anytime guarantees on the agent’s performance, assuming smoothness in the MDP. However, it is possible to replace this assumption with a weaker one, at the cost of only retaining guarantees in expectation. This weaker version of the theorem can be retrieved by simply assuming smoothness on the noise, rather than on the MDP, and leveraging results recently presented by Maran et al. (2023).
Assumption 5.
The noise distribution $\bm{\epsilon}$ is $L_{\ell}$-TV-Lipschitz continuous.
This assumption is satisfied by a large class of Gaussian and sub-Gaussian distributions (Maran et al., 2023). We can build upon Assumption 4 and Lemma 2, and start by providing a weaker version of Lemma 3.
Lemma 6.
(Uniform convergence of marginal variance in expectation) Under Assumption 1, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
$\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\sigma^{2}_{n}(s,c)\leq\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau^{2}_{\text{min}}},$
where $\tilde{\sigma}^{2}=\max_{(s,c)\in\mathcal{S\times C}}\sigma^{2}_{0}(s,c)+\rho^{2}(s,c)$ and $\tau_{\text{min}}=\min_{s,c\in\mathcal{S\times C}}\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(c)}\mathbf{1}_{s\in\tau}$.
Proof.
We resume from Inequality 60 in the proof of Lemma 3:
$\sigma_{n}^{2}(s,c)\leq\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (79)
Therefore,
$\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\sigma^{2}_{n}(s,c)$ $\leq\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau\sim\bm{\tau}(c)\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (80)
$\leq\max_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau^{\prime},c^{\prime})\mid c_{1:n},\tau_{1:n})$ (81)
$=\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau^{2}_{\text{min}}}.$ (82)
Having bounded variance at each round, this time in expectation, we can invoke Lemma 4 to bound the expected distance to the optimal policy with high probability:
$\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\|\pi^{\star}(s,c)-\mu_{n}(s,c)\|_{1}\leq\beta^{\prime}_{n}(\delta)\|\bar{\sigma}\|_{1}\Big{(}\frac{2\Gamma_{n}}{\tau^{2}_{\text{min}}}\Big{)}^{\frac{1}{2}}$ (83)
Instead of leveraging bounds for the imitator’s performance in Lipschitz-smooth settings, we can instead use the fact that the expert’s actions are corrupted by smooth noise. In this setting, it is instead possible to control the suboptimality of the imitator with respect to the noisy expert. We report the following Theorem from Maran et al. (2023), and refer to the original work for the proof.
Lemma 7.
Let $\pi^{\star}$, $\bm{\tilde{\pi}}$ and $\pi$ denote the expert, noisy expert and imitator policy, respectively. If Assumption 5 holds, then:
$J^{\bm{\tilde{\pi}}}-J^{\pi}\leq\frac{2L_{\ell}Q_{\text{max}}}{1-\gamma}\mathbb{E}_{s\sim\mu_{\bm{\tilde{\pi}}}}[\mathcal{W}(\pi^{\star}(\cdot\mid s),\pi(\cdot\mid s))],$
where $Q_{\text{max}}=\max_{(s,c,a)\in\mathcal{S}\times\mathcal{C}\times\mathcal{A}}|Q(s,a)|^{\pi}$ and $\mathcal{W}$ represents the Wasserstein $1$-distance.
As the expert $\pi^{\star}$ and the imitator $\pi_{n}$ are both deterministic, this implies that
$J^{\bm{\tilde{\pi}}}-J^{\pi}_{n}\leq\frac{2L_{\ell}Q_{\text{max}}}{1-\gamma}\mathbb{E}_{s\sim\mu_{\bm{\tilde{\pi}}}}\|(\pi^{\star}(\cdot\mid s),\pi(\cdot\mid s))\|_{1}.$ (84)
By invoking this Lemma, we can thus conclude that, with probability at least $1-\delta$
$\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}J^{\bm{\tilde{\pi}}}-J^{\pi_{n}}\leq\frac{2^{\frac{3}{2}}L_{\ell}Q_{\text{max}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\beta^{\prime}_{n}(\delta)\Gamma_{n}^{\frac{1}{2}}=O(\gamma_{(Hn)}n^{-\frac{1}{2}}).$ (85)
Therefore, if $\gamma_{n}=O(\log n)$, then $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}J^{\bm{\tilde{\pi}}}-J^{\pi}=O(\frac{\log n}{\sqrt{n}})\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$. While these performance guarantees only hold in expectation, they arise from minimal assumptions, mostly regarding the policy class and the perturbation noise, and can thus be applied to arbitrary MDPs.
Appendix C Practical Objective
Following up on the approximations reported in Section 4.2, we present the empirical estimate of the objective that is used through experiments. In particular, we show how the expectations in Equation 2 may be approximated with finite samples. The original criterion is expressed as
$c_{n}=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\phi_{n}(c^{\prime})=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}).$ (86)
An empirical estimate can be derived as follows:
$\phi_{n}(c^{\prime})$ $=\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$ (87)
$\stackrel{{\scriptstyle(i)}}{{\approx}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\hat{\tau}_{1:n-1})$ (88)
$\stackrel{{\scriptstyle(ii)}}{{\approx}}\frac{1}{|\mathcal{\hat{C}}|}\sum_{c\in\mathcal{\hat{C}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$ (89)
$=\frac{1}{|\mathcal{\hat{C}}|}\sum_{c\in\mathcal{\hat{C}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau^{\prime}\sim\bm{\hat{\tau}}\\ (s_{0},\dots)\sim\bm{\hat{\tau}}\end{subarray}}\frac{\bm{\tau}(\tau^{\prime}|c^{\prime})}{\bm{\hat{\tau}}(\tau^{\prime})}\frac{\bm{\tau}((s_{0},\dots)|c)}{\bm{\hat{\tau}}((s_{0},\dots))}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$ (90)
$=\frac{1}{|\mathcal{\hat{C}}|}\sum_{c\in\mathcal{\hat{C}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau^{\prime}\sim\bm{\hat{\tau}}\\ (s_{0},\dots)\sim\bm{\hat{\tau}}\end{subarray}}w(\tau^{\prime},c^{\prime})w((s_{0},\dots),c)\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$ (91)
$\stackrel{{\scriptstyle(iii)}}{{\approx}}\frac{1}{|\mathcal{\hat{C}}|(n-1)^{2}}\sum_{c\in\mathcal{\hat{C}}}\sum_{\begin{subarray}{c}\tau^{\prime}\in\hat{\tau}_{1:n-1}\\ (s_{0},\dots)\in\hat{\tau}_{1:n-1}\end{subarray}}w(\tau^{\prime},c^{\prime})w((s_{0},\dots),c)\sum_{t=0}^{H-1}\!\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$ (92)
where (i) uses a single sample to estimate the expectation over past trajectories, (ii) uses a sample-based approximation to the target task distribution $\mu_{c}$, and (iii) uses the importance sampling trick introduced in Section 4.2, with $w(\tau,c)=\frac{(n-1)\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c)}{\sum_{i=0}^{n-1}\prod_{t=0}^{H-1}\bm{\tilde{\pi}}(a_{t}|s_{t},c_{i})}$. This final approximate objective does not involve expectations, and can be efficiently computed. The complexity of evaluating the criterion for a single task $c^{\prime}$ scales linearly with the number of samples in $\mathcal{\hat{C}}$ and quadratically with the number of rounds $n$. However, the dependency on the number of rounds can be removed by evaluating the second sum over a fixed number of trajectories sampled among $\tau_{1:n-1}$, ensuring that the complexity does not depend on the round.
Appendix D Additional results for AMF-GP
Figure 3 only reports full return curves for two representative pre-training settings, namely those involving $6/12$ and $/12$ 12 demonstrated tasks. We here report full results for each task allocation, spanning from $1/12$ to $/12$ 12 demonstrated tasks. For each setting, we report both average multi-task return and average policy entropy curves.
1/12 tasks
2/12 tasks
3/12 tasks
4/12 tasks
5/12 tasks
6/12 tasks
7/12 tasks
8/12 tasks
9/12 tasks
10/12 tasks
11/12 tasks
12/12 tasks
  AMF-GP   Uniform
Figure 8: Additional results in GP settings for a 2D integrator (see Figure 2). AMF-GP results in improved sample efficiency across all pre-training regimes, and is particularly effective for skewed pre-training distributions (e.g., when pre-training demonstrations have been allocated to 1/12 or 6/12 tasks).
Appendix E Additional results for AMF-NN
Results in Figure 4 are computed over two representative pre-training distributions: one allocating pre-training demonstrations uniformly over all tasks, the other one only demonstrating the first two tasks. We report these results again, and compare them with those for several other pre-training distributions. In particular, we evaluate a family of skewed priors which are only trained on one or two tasks. Results are reported for FrankaKitchen in Figure 9 and for Metaworld in Figure 10, and are consistent with patterns observed in Figure 4.
0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
$[3,3,3,3,3]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[8,8,0,0,0]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,8,8,0,0]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,0,8,8,0]$
0510Rounds0.20.40.60.8Multi-task success rateKitchen, skewed
$[0,0,0,8,8]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[16,0,0,0,0]$
  AMF-NN ($\sigma=1e^{-3}$)    Uniform    Uniform, no parameter isolation
Figure 9: Additional results for AMF-NN in FrankaKitchen with state inputs. We evaluate several allocations of the pre-training demonstrations, as labeled below each plot (e.g., the label $[8,8,0,0,0]$ indicates that 8 demonstrations were provided for each of the first two tasks each, and none for the remaining tasks).
0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
$[4,4,4,4]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[8,8,0,0]$
0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
$[0,8,8,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[0,0,8,8]$
0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
$[16,0,0,0]$
  AMF-NN ($\sigma=1e^{-3}$)   Uniform    Uniform, no parameter isolation
Figure 10: Additional results for AMF-NN in Metaworld with state inputs. We evaluate several allocations of the pre-training demonstrations, as labeled below each plot (e.g., the label $[8,8,0,0]$ indicates that 8 demonstrations were provided for each of the first two tasks each, and none for the remaining tasks).
Moreover, we also report an ablation over the choice of batch size for Metaworld in Figure 11, thus complementing the one reported in Section 5.3. We observe a slight upward trend favoring larger batch size in the case of an uniform prior, but smaller batches remain overall desirable.
34567Area under success rate curveMetaworld, uniform
34567Area under success rate curveMetaworld, skewed
$B\!=1$
$B\!=2$
$B\!=4$
Uniform
Uniform, no
parameter isolation
Figure 11: Influence of batch size $B$ over area under success rate curve with a budget of 20 demonstrations in Metaworld.
Appendix F Uncertainty ablation
Section 5.4 evaluates alternative uncertainty quantification schemes in FrankaKitchen for two pre-training distributions. This Section extends these results to include results for Metaworld, and for several other pre-training settings (see Figures 12 and 13). Results are consistent with those so far reported, suggesting that loss gradient embeddings are an important component for the empirical performance of AMF-NN.
0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
$[3,3,3,3,3]$
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
$[8,8,0,0,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,8,8,0,0]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,0,8,8,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,0,0,8,8]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[16,0,0,0,0]$
  Last-layer gradient    Last-layer    Dropout
Figure 12: Additional results for AMF-NN in FrankaKitchen with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
0510Rounds0.550.600.650.700.75Multi-task success rateMetaworld, uniform
$[4,4,4,4]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[8,8,0,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[0,8,8,0]$
0510Rounds0.50.60.70.8Multi-task success rateMetaworld, skewed
$[0,0,8,8]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[16,0,0,0]$
  Last-layer gradient    Last-layer    Dropout
Figure 13: Additional results for AMF-NN in Metaworld with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
Appendix G Mitigating forgetting
The ability of neural networks to adapt to shifts in training distribution while retaining information is an important object of interest in lifelong and continual learning (Wang et al., 2024). In general, learned models display a trade-off between their ability to integrate novel information, and their memory of previously observed training samples. Arguably, common neural network architectures can easily fit new data (save for loss of plasticity (Lyle et al., 2023)), but are known to forget previous information, often catastrophically. This problem is of utmost relevance in our setting, in which the pre-trained network is not just leveraged as a useful initialization, but may already capable of solving some tasks. Hence, the fine-tuning procedure should be careful not to disrupt this ability.
Several methods aimed at mitigating forgetting can be traced back to rehearsal (Riemer et al., 2019; Chaudhry et al., 2019) and regularization (Kirkpatrick et al., 2017) strategies. While rehearsal approaches are often effective, they also require access to pre-training data, which is unrealistic in our setting. Hence we consider two common regularization technique, namely L2-regularization to the pre-trained weights, and EWC (Kirkpatrick et al., 2017). The latter can be seen as a more nuanced version of the former, which adaptively scales the regularization strength according to the curvature of the loss landscape. Unfortunately, we found both methods to be insufficient in our setting, as reported in Figure 14. When coupled with large regularization weights, their asymptotical performance is significantly limited. When regularization weights are too low, they recover the performance of a naive baseline. Intermediate values were found to interpolate between the two behaviors, without addressing the forgetting issue. While less scalable, we found hard parameter isolation to be the only effective solution amongst the one we tested. As mitigating forgetting in neural network is an orthogonal direction to the main topic of this work, we adopt this solution, and expect that further developments in continual learning will be applicable in our setting.
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
0510Rounds0.40.50.60.7Multi-task success rateMetaworld, uniform
0510Rounds0.20.40.6Multi-task success rateMetaworld, skewed
  Parameter isolation    EWC    L2 regularization    Baseline
Figure 14: Performance of uniform task sampling with several techniques to mitigate forgetting. Darker shades represent stronger regularization coefficients. L2 and EWC regularization are not effective in this setting.
Appendix H Criterion vs returns
0.00.10.20.3Pretraining distribution0.00.51.01.51e-4Estimated expected posterior entropyTask space-13.5-13.0-12.5Average return after finetuning
Figure 15: Didactic example on correlation between pre-training distribution over tasks (top), evaluations of the AMF criterion for each task (middle) and return after fine-tuning on a demonstration for a given task (bottom).
As a didactic example, we evaluate the criterion optimized by AMF-GP in a particular instance. We adopt the settings presented in Section 5.1, and pre-train a GP policy by providing 50 demonstrations in the 2D integrator environment, uniformly sampled among tasks in the top half of the target circle. We represent the task space along one dimension, and plot the smoothed pre-training distribution on the top of Figure 15. The second row of the Figure displays the evaluation of the criterion in Equation 2 for 100 tasks uniformly sampled across the entire task space. By comparison with the plot above, it is evident that the criterion is significantly lower for tasks that have not yet been demonstrated. These tasks are also those that, if demonstrated, would lead to a greater increase in multi-task performance after fine-tuning, as reported in the bottom row of Figure 15. In this instance, it’s easy to see that the criterion leads to selection of tasks which have not been demonstrated sufficiently, and that will thus lead to greater policy performance.
Appendix I Implementation Details
In order to ease reproducibility, we open-source our codebase on the project’s anonymous website.444sites.google.com/active-multitask-finetuning Furthermore, we describe several implementation details in the following sections.
I.1 Metrics
All metrics are reported in the form of their mean and the 90% simple bootstrap confidence intervals over $10$ random seeds.
I.2 GP settings
In GP settings (5.1), each expert demonstration involves $5$ steps, is corrupted with Gaussian noise and collected by a scripted policy. As the task space is continuous, the criterion is simply optimized via uniform random shooting, with a budget of $100$. Multi-task returns are averaged over $20$ episodes per task.
I.3 Neural network settings
I.3.1 Environments
We evaluate AMF-NN across three environments, namely FrankaKitchen, Metaworld and WidowX. For the first two, demonstrations are $50$ steps, while for the latter they involve $100$ steps. In FrankaKitchen, demonstrations are provided by Kumar et al. (2024), and collected by trained SAC agents. In Metaworld, demostrations are instead collected by the scripted policies provided (Yu et al., 2020). Finally, in WidowX successful trajectories are collected by Octo-small (Octo Model Team et al., 2024) itself and filtered according to success labels, in an instance of self-supervised distillation. Furthermore, in the case of WidowX, the initial position of the object is not randomized, as we found this to result in very inconsistent performance for the data collection policy. In the first two environments, $25$ attempts for each task are evaluated, while the evaluation for WidowX involves $50$ attempts.
I.3.2 Behavior cloning
The policy is parameterized via a deterministic MLP with $2$ layers and $256$ units per layer, with layer normalization (Ba, 2016). Task conditioning are image embedding extracted by R3M (Nair et al., 2022a). Policies are pre-trained for $500$ epochs with batch size of $128$, learning rate of $41.e-$ 4 using the Adam optimizer (Kingma, 2014).
I.3.3 Importance sampling weights
In GP setting, importance weights are computed from the Gaussian policy distribution, and log-probabilities are clipped to the range $[-12,0]$. In NN settings, for deterministic policies, we interpret the policy’s output as the mean of a Gaussian with fixed standard deviation $\sigma=1.0$, and only clip log-probabilities for numerical stability. In experiments involving pre-trained Octo policies, we evaluated two solutions. One option consisted of fitting a Gaussian distribution through maximum likelihood methods to samples from the diffusion policy, and was found to underperform. We thus treat the Octo policy as strictly deterministic: with continuous action spaces, this simplifies importance sampling weights to $w(c,\tau)=1$ in case $\tau$ is a demonstration provided exactly for task $c$, and 00 otherwise. We note that this solution cannot be used to evaluate the criterion on yet unobserved tasks, but remains feasible when tasks are finite and few.
I.3.4 AMF
Each fine-tuning round involves $3000$ gradient steps, each with a batch size of $128$. We warm-start each algorithm by collecting the first $|\mathcal{C}|$ demonstrations uniformly, as mentioned in Section 4.2. In the case of loss-gradient embeddings, we found it to be beneficial to use a separate copy of the policy for task selection, which is not trained on these initial trajectories (which thus can be seen as a small “validation” set). As these demonstrations are not selected according to the criterion, to avoid unwanted updates of pre-training weights, they are only added to the training set once the algorithm selects a task belonging to the same partition of the task space, as defined by parameter isolation.
Appendix J Useful inequalities
Lemma 8.
If $a,b\in(0,M]$ for some $M>0$ and $b\geq a$ then
$b-a\leq M\cdot\log\left(\frac{b}{a}\right).$
If additionally, $a\geq M^{\prime}$ for some $M^{\prime}>0$ then
$b-a\geq M^{\prime}\cdot\log\left(\frac{b}{a}\right).$
Proof.
[…]
$\frac{1}{c}=f^{\prime}(c)=\frac{f(b)-f(a)}{b-a}=\frac{\log b-\log a}{b-a}=\frac{\log\left(\frac{b}{a}\right)}{b-a}.$
Thus,
$b-a=c\cdot\log\left(\frac{b}{a}\right)<M\cdot\log\left(\frac{b}{a}\right).$
Under the additional condition that $a\geq M^{\prime}$, we obtain
$b-a=c\cdot\log\left(\frac{b}{a}\right)>M^{\prime}\cdot\log\left(\frac{b}{a}\right).$
Lemma 9.
Let us consider two spaces $X\in\mathbb{R}^{n}$, $Y\in\mathbb{R}^{m}$, and a conditional distribution $p:X\to\Delta(Y)$ whose support $\text{supp}(p(\cdot|x))$ is bounded by a ball of radius $\epsilon$ for all $x\in X$, that is
$\max_{y_{l},y_{h}\in\text{supp}p(\cdot|x)}\|y_{h}-y_{l}\|_{2}\leq\epsilon.$
For all $(x,x^{\prime})\subseteq X$, $y\sim p(\cdot|x)$ and $y^{\prime}\sim p(\cdot|x^{\prime})$ it holds that
$\|y-y^{\prime}\|_{2}\leq\mathcal{W}(p(\cdot|x),p(\cdot|x^{\prime}))+2\epsilon,$
where $K$ denotes the Wasserstein $1$-distance.
Proof.
$\|y-y^{\prime}\|$ $\leq\max_{\begin{subarray}{c}y\in\text{supp}(p(\cdot|x))\\ y^{\prime}\in\text{supp}(p(\cdot|x^{\prime}))\end{subarray}}\|y-y^{\prime}\|_{2}$ (93)
$\stackrel{{\scriptstyle(i)}}{{\leq}}\min_{\begin{subarray}{c}y\in\text{supp}(p(\cdot|x))\\ y^{\prime}\in\text{supp}(p(\cdot|x^{\prime}))\end{subarray}}\|y-y^{\prime}\|_{2}+2\epsilon$ (94)
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathcal{W}(p(\cdot|x),p(\cdot|x^{\prime}))+2\epsilon,$ (95)
where (i) follows from the triangle inequality, and (2) is due to the fact that the integral of the distance between two points in $Y$ for any coupling is greater than the minimum distance. ∎

### v3

@@
-Active Fine-Tuning of Generalist Policies
+Active Fine-Tuning of Multi-Task Policies
 Abstract
@@
 We derive performance guarantees for AMF under regularity assumptions and demonstrate its empirical effectiveness to efficiently fine-tune neural policies in complex and high-dimensional environments.
+Imitation Learning, Multi-task Learning, Active Learning, Deep Reinforcement Learning
 1 Introduction
+The availability of large pre-trained models has transformed entire areas of machine learning, from computer vision (Krizhevsky et al., 2012; He et al., 2016; Dosovitskiy et al., 2021; Radford et al., 2021), to natural language processing (Radford et al., 2019; Brown et al., 2020) and generative modeling in general (Ho et al., 2020; Esser et al., 2024).
+This paradigm has started to extend to robotics and control (Collaboration, 2023; Ma et al., 2024), in particular for systems for which demonstrations are readily available (Octo Model Team et al., 2024), or can be easily collected (Zhao et al., 2023).
+Even when demonstrations are not easily obtained, scaling laws in reinforcement learning (Ceron et al., 2024b, a; Nauman et al., 2024) suggest the possibility of leveraging large pre-trained policies.
+These “generalist” policies have decent performance on many tasks, and can be fine-tuned on particular set of tasks while leveraging their previously learned representations and skills.
+We investigate whether representations of pre-trained policies can be used to significantly bootstrap learning progress.
 Figure 1: Interactive loop between agent and expert.
@@
 We study how to select tasks (in blue) to obtain the best-performing policy after as few demonstrations as possible.
-The availability of large pre-trained models has transformed entire areas of machine learning, from computer vision (Krizhevsky et al., 2012; He et al., 2016; Dosovitskiy et al., 2021; Radford et al., 2021), to natural language processing (Radford et al., 2019; Brown et al., 2020) and generative modeling in general (Ho et al., 2020; Esser et al., 2024).
-This paradigm has started to extend to robotics and control (Collaboration, 2023; Ma et al., 2024), in particular for systems for which demonstrations are readily available (Octo Model Team et al., 2024), or can be easily collected (Zhao et al., 2023).
-Even when demonstrations are not easily obtained, scaling laws in reinforcement learning (Ceron et al., 2024b; a; Nauman et al., 2024) suggest the possibility of leveraging large pre-trained policies.
-These “generalist” policies have decent performance on many tasks, and can be fine-tuned on particular set of tasks while leveraging their previously learned representations and skills.
-We investigate whether representations of such policies can be used to significantly bootstrap learning progress.
-As a motivating example, consider a household robot that is delivered with a pre-trained “generalist” policy, and deployed in slightly different conditions than those observed in its training data.
+As a motivating example, consider a household robot that is delivered with a pre-trained “generalist” policy, and deployed in different conditions than those observed in its training data.
 While the robot may achieve some tasks in a zero-shot fashion (e.g.
@@
 AMF parallels recent work on active supervised fine-tuning of neural networks (Hübotter et al., 2024).
-To this end, AMF relies on estimates of the demonstrations’ information gain about the expert policy.
+To this end, AMF relies on estimates of the demonstrations’ information gain on the expert policy.
 We prove that in sufficiently regular Markov decision processes, AMF converges to the expert policy.
@@
   • We empirically scale AMF to high-dimensional tasks involving pre-trained neural policies.
+  • We additionally propose a practical approach to alleviate catastrophic forgetting in pre-trained neural policies, which benefits arbitrary data selection strategies.
 2 Related Work
 Learning-based control and active data selection are both well-established research directions.
-This section discusses some of the most topical works in either direction, and clarifies the novelty and placement of this work with respect to them.
-Behavioral Cloning
-Numerous imitation learning approaches have been developed with the goal of distilling knowledge from high-quality demonstrations to a control policy (Osa et al., 2018).
+This section discusses some topical works in either direction, and clarifies the novelty and placement of this work with respect to them.
+Behavioral Cloning  Numerous imitation learning approaches have been developed with the goal of distilling knowledge from high-quality demonstrations to a control policy (Osa et al., 2018).
 Within this family of techniques, behavioral cloning (BC, Bain & Sammut, 1995; Ross & Bagnell, 2010) aims to maximize policy performance by minimizing the distance of its actions to demonstrated actions, simply through supervised learning.
@@
 Next to recent empirical successes (Chi et al., 2023), formal analysis has also advanced (Spencer et al., 2021; Block et al., 2024a; Belkhale et al., 2024; Foster et al., 2024), and established provable performance guarantees for BC policies (Xu et al., 2020; Maran et al., 2023; Block et al., 2024b).
-Multi-task and Generalist Policies
-Traditionally, behavioral cloning has mostly been deployed in a single-task setting.
+Multi-task and Generalist Policies  Traditionally, behavioral cloning has mostly been deployed in a single-task setting.
 Multi-task learning in sequential decision-making has largely been investigated in the context of reinforcement learning (Teh et al., 2017; Sodhani et al., 2021; Yu et al., 2021; Sun et al., 2022; Cho et al., 2022; Hendawy et al., 2023).
@@
 We find this capability to emerge even from models that are not explicitly trained in this way, and focus on which demonstrations to obtain.
-Data Selection
-The idea of directing a sampling process to gather information has been central to machine learning research and studied extensively in experimental design (Chaloner & Verdinelli, 1995) and active learning (Settles, 2009).
-Most work on active data selection summarizes data without focusing on a particular task (e.g., Sener & Savarese, 2017; Ash et al., 2020b; Holzmüller et al., 2023; Lightman et al., 2023), which has been predominantly applied to pre-training.
+Data Selection  The idea of directing a sampling process to gather information has been central to machine learning research and studied extensively in experimental design (Chaloner & Verdinelli, 1995), active learning (Settles, 2009) and reinforcement learning (Mehta et al., 2022).
+Most work on active data selection summarizes data without focusing on a particular task (e.g., Sener & Savarese, 2017; Ash et al., 2020; Holzmüller et al., 2023; Lightman et al., 2023), which has been predominantly applied to pre-training.
 Recently, adapting models after pre-training and during deployment has gained interest.
@@
 We focus instead on supervised fine-tuning of learning-based controllers in dynamical systems.
-This necessitates automatic data selection, for which practical methods currently rely on uniform sampling or externally provided heuristics.
-Our approach extends work on task-directed data selection (Kothawade et al., 2020; Wang et al., 2021a; Kothawade et al., 2022; Smith et al., 2023), which has recently been applied to the supervised fine-tuning of large-scale neural networks in vision (Hübotter et al., 2024) and language (Xia et al., 2024).
+Our approach extends work on task-directed data selection (Kothawade et al., 2020; Wang et al., 2021a; Kothawade et al., 2022; Bickford Smith et al., 2023), which has recently been applied to the supervised fine-tuning of large-scale neural networks in vision (Hübotter et al., 2024) and language (Rotman & Reichart, 2022; Xia et al., 2024; Hübotter et al., 2025).
 3 Background
@@
 Any given policy induces a task-conditional distribution over trajectories:
-$\bm{\tau}_{\bm{\pi}}\Big{(}\big{(}s_{0},a_{0},s_{1},a_{1},\dots\big{)}\mid c\Big{)}=\mu_{0}(s_{0})\prod_{t=0}^{\infty}\bm{\pi}(a_{t}\mid s_{t},c)\cdot P(s_{t+1}\mid s_{t},a_{t}).$
+$\bm{\tau}_{\bm{\pi}}$ $\Big{(}\big{(}s_{0},a_{0},s_{1},a_{1},\dots\big{)}\mid c\Big{)}=$
+$\mu_{0}(s_{0})\prod_{t=0}^{\infty}\bm{\pi}(a_{t}\mid s_{t},c)\cdot P(s_{t+1}\mid s_{t},a_{t}).$
 The discounted returns for a specific task $c\in\mathcal{C}$ or a task distribution $\mu_{c}\in\Delta(\mathcal{C})$ are, respectively,
@@
 In the scope of this work, we instead consider an imitation learning setting, in which expert demonstrations from an optimal policy $\pi^{\star}$ are provided.
-In particular, we focus on behavioral cloning algorithms, which reduce the control problem to a supervised learning problem.
-Given a set of $N$ task-conditioned, $H$-length trajectories $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ with task labels $c_{1:N}$, behavioral cloning proposes a proxy objective for the policy $\pi$: an empirical estimate of the log-likelihood under the data distribution: $J^{\bm{\pi}}_{\text{proxy}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\log\bm{\pi}(a_{t}^{i}\mid s_{t}^{i},c_{i}).$ If trajectories $\tau_{1:N}$ are obtained from the optimal policy, cover the support of the desired task distribution $\mu_{c}$, and the searched policy class is sufficiently rich, the maximizer of $J^{\bm{\pi}}_{\text{proxy}}$ will also maximize $J^{\bm{\pi}}_{\mu_{c}}$ as $N$ and $H$ increase.
+In particular, we focus on behavioral cloning algorithms, which reduce control to a supervised learning problem.
+Given a set of $N$ task-conditioned, $H$-length trajectories $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ with task labels $c_{1:N}$, behavioral cloning proposes a proxy objective for the policy $\bm{\pi}$: an empirical estimate of the log-likelihood under the data distribution: $J^{\bm{\pi}}_{\text{proxy}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\log\bm{\pi}(a_{t}^{i}\mid s_{t}^{i},c_{i}).$ If trajectories $\tau_{1:N}$ are obtained from the optimal policy, cover the support of the desired task distribution $\mu_{c}$, and the searched policy class is sufficiently rich, the maximizer of $J^{\bm{\pi}}_{\text{proxy}}$ will also maximize $J^{\bm{\pi}}_{\mu_{c}}$ as $N$ and $H$ increase.
 However, in general, there is a clear mismatch between $J^{\bm{\pi}}_{\text{proxy}}$ and $J^{\bm{\pi}}_{\mu_{c}}$ (Xu et al., 2020; Maran et al., 2023).
@@
 The process is repeated for $N$ rounds, with the goal of producing a fine-tuned policy that maximizes the expected returns for the desired task distribution $\mu_{c}$.
-Modeling assumptions
-We take a Bayesian perspective on active multi-task fine-tuning, by assuming a Bayesian model $\bm{\pi}$ over policies.
+We note that the fine-tuning process does not assume access to the pre-training data distribution.
+This setting is both realistic, as large pre-training datasets are rarely publicly available, and challenging, as it prohibits any naive data rebalancing strategy.
+Modeling assumptions   We take a Bayesian perspective on active multi-task fine-tuning, by assuming a Bayesian model $\bm{\pi}$ over policies.
 We assume that demonstrations follow a noisy expert: $\bm{\tilde{\pi}}(s,c)=\pi^{\star}(s,c)+\bm{\epsilon}(s,c)$ where $\bm{\epsilon}(s,c)$ is independent noise.
@@
 We thus build on top of principled active learning approaches for non-sequential domains (Hübotter et al., 2024), and propose AMF, which selects queries that maximize the expected information gain about the expert policy over its occupancy:
-$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tau}(c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}).$ (1)
+$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tau}(c^{\prime})\mid c_{1:n-1},\tau_{1:n-1}),$
+$\text{with }c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c),\;\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}).$ (1)
 We show in Section 4.1 that, under certain regularity assumptions, the policy learned by AMF converges to the expert policy and matches its performance.
 These results constitute a first-of-its-kind performance guarantee for active multi-task fine-tuning.
-The main novelty of this guarantee is the extension of prior work to sequential domains where the visited trajectory $(s_{0},a_{0},s_{1},\dots)\sim\bm{\tau}(c_{n})$ is unknown when selecting the task $c_{n}$ for a demonstration.
+The main novelties of this guarantee are the extension of prior work to sequential domains where the visited trajectory $(s_{0},a_{0},s_{1},\dots)\sim\bm{\tau}(c_{n})$ is unknown when selecting the task $c_{n}$ for a demonstration, and the connection to imitator performance.
 In Section 4.2, we discuss the design choices that make AMF amenable to optimization in practical settings.
@@
 We summarize the main result here, and include a formal proof in Appendix A.
-Informal Assumption 1.
-We make the following assumptions:
+Informal Assumption 4.1.
+We make these assumptions:
   1. 1.
@@
 Under these assumptions, we prove the following performance guarantee for active multi-task behavioral cloning.
-Informal Theorem 2 (Performance guarantees for active multi-task BC).
+Informal Theorem 4.2 (Performance guarantees for active multi-task BC).
 Let all regularity assumptions hold.
 If each demonstrated task of length $H$ is selected according to the criterion in Equation 1, then with probability $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
-$J^{\pi^{\star}}_{\mu_{c}}-J^{\pi_{n}}_{\mu_{c}}\leq O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
+$J^{\pi^{\star}}_{\mu_{c}}-J^{\pi_{n}}_{\mu_{c}}\leq O(\gamma_{(Hn)})/\sqrt{n},$
 where $\pi_{n}$ is the mean of $\bm{\pi}$ at round $n$ and $\gamma_{(Hn)}$ is the maximum information gain about the expert policy from $Hn$ samples, and is sublinear for a large class of problems.
 The $O(\cdot)$ notation suppresses all multiplicative terms that do not depend on $n$.
-Intuitively, this theorem proves that the imitator will eventually achieve the demonstrator’s performance in smooth, regular MDPs with sublinear $\gamma_{(Hn)}$ (for a formal definition, we refer to Lemma 2 in the Appendix).
+Intuitively, this theorem proves that the imitator will eventually achieve the demonstrator’s performance in smooth, regular MDPs with sublinear $\gamma_{(Hn)}$ (for a formal definition, we refer to Lemma A.6 in the Appendix).
 We can also prove a more general result under weaker assumptions: as long as the policy is regular, the imitator will reach the noisy expert performance in arbitrary, non-smooth MDPs, albeit only in expectation.
@@
 4.2 Practical Algorithms
-Theorem 2 guarantees that, under regularity assumptions, the adaptive demonstration sampling scheme leads to convergence of the imitator’s performance to the optimal one.
+Theorem 4.2 guarantees that, under regularity assumptions, the adaptive demonstration sampling scheme leads to convergence of the imitator’s performance to the optimal one.
 However, this criterion involves state occupancies and a conditional entropy term, which are hard to access or estimate in practice.
@@
 We first rephrase the objective from Equation 1 in its entropy form:
-$c_{n}$ $=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$ (2)
+$c_{n}=$ $\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1}),$
+with $c\sim\mu_{c},(s_{0},\dots)\sim\bm{\tau}(c),\tau^{\prime}\sim\bm{\tau}(c^{\prime}),$
+$\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}).$ (2)
 where we use the definition of mutual information $\mathcal{I}(\cdot|\bm{\tau}(c^{\prime}))=\mathcal{H}(\cdot)-\mathcal{H}(\cdot|\bm{\tau}(c^{\prime}))$, drop the first entropy term as it does not depend on $c^{\prime}$, and rewrite the second entropy term as an expectation over $\bm{\tau}(c^{\prime})$.
@@
 However, the remaining two expectations (i.e., $\mathop{\mathbb{E}}_{\tau^{\prime}\sim\bm{\tau}(c^{\prime})}(\cdot)$ and $\mathop{\mathbb{E}}_{(s_{0},\dots)\sim\bm{\tau}(c)}(\cdot)$) involve the distribution over trajectories for an arbitrary task, which might not have been demonstrated yet.
-However, we observe that, at round $n$, the tasks demonstrated so far induce the empirical distribution $\hat{\mu}_{c}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{c_{i}}(\cdot)$, while the trajectories collected similarly induce $\bm{\hat{\tau}}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{\hat{\tau}_{i}}(\cdot)$, where $\delta$ indicates the Dirac delta distribution.
+We observe that, at round $n$, the tasks demonstrated so far induce the empirical distribution ${\hat{\mu}_{c}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{c_{i}}(\cdot)}$, while the trajectories collected similarly induce ${\bm{\hat{\tau}}(\cdot)=\frac{1}{n-1}\sum_{i=1}^{n-1}\delta_{\hat{\tau}_{i}}(\cdot)}$, where $\delta$ indicates the Dirac delta distribution.
 We can show that expectations over the trajectory distribution for an arbitrary task $c\in\mathcal{C}$ can be estimated through importance sampling (i.e., by sampling trajectories from $\bm{\hat{\tau}}(\cdot)$ instead of $\bm{\tau}(\cdot|c)$): ${\mathop{\mathbb{E}}_{\tau\sim\bm{\tau}(\cdot|c)}f(\tau)=\mathop{\mathbb{E}}_{\tau\sim\bm{\hat{\tau}}(\cdot)}\frac{\bm{\tau}(\tau|c)}{\bm{\hat{\tau}}(\tau)}f(\tau)}$.
@@
 On the other hand, the high-variance of the estimate can be controlled by practical solutions such as clipping.
-We present the resulting empirical estimate for Equation 2 in full in Appendix C.
+We present the resulting empirical estimate for Equation 2 in full in Appendix C, and a qualitative analysis of importance weights in Appendix K.
 Entropy estimation
@@
 The unconditional entropy can be measured in closed form as ${\mathcal{H}(\bm{\pi}(x))=\frac{1}{2}\log(2\pi k(x,x))+\frac{1}{2}}$, and the conditional entropy can be obtained by simply replacing the kernel $k$ with $\hat{k}_{X}(x,x)=k(x,x)-k(x,X)[k(X,X)+\sigma_{\epsilon}^{2}I]^{-1}k(X,x),$ where $\sigma_{\epsilon}^{2}$ is the variance of the observation noise $\bm{\epsilon}(s,c)$, assuming it is distributed according to a zero-mean Gaussian.
-Algorithm 1 Practical AMF
+Algorithm 1 AMF
   Input: initial policy $\pi_{0}$, budget $N$, desired task distr. $\mu_{c}$
@@
   for $n\in[0,\dots,N-1]$ do
-     Compute $c_{n}$ as the solution to Equation 2
+     Compute $c_{n}$ as the solution to Eq. 2
      Collect new demonstration $\tau_{n}$ for task $c_{n}$
-     Update $\pi_{n+1}$ from $\pi_{n}$ with $\mathcal{D}_{n}=\mathcal{D}_{n-1}\cup\{c_{n},\tau_{n}\}$
+     if $n+1\;\%\;B=0$ then
+        $\mathcal{D}_{n+1}=\mathcal{D}_{n+1-B}\cup\{c_{n-B+1:n},\tau_{n-B+1:n}\}$
+        Update $\pi_{n+1}$ from $\pi_{n+1-B}$ with $\mathcal{D}_{n+1}$
+     end if
   end for
@@
 Application of the method to policies parameterized by neural networks will adopt the same scheme.
-It will also require additional care on three distinct topics: (i) kernel approximations, (ii) batch selection and (iii) forgetting mitigation.
+It will also require additional care on two distinct topics: kernel approximtions and mitigation of catastrophic forgetting.
 Kernel approximations
@@
 This technique does not violate any of the practical constraints listed above, and allows us to adapt the machinery introduced in GP settings.
-While several embedding strategies exist Jacot et al.
-(2018); Devlin et al.
-(2019); Holzmüller et al.
-(2023), we adopt loss gradient embeddings (Ash et al., 2020a).
+While several embedding strategies exist (Jacot et al., 2018; Devlin et al., 2019; Holzmüller et al., 2023), we adopt loss gradient embeddings (Ash et al., 2020).
 Assuming the prior $\bm{\beta}\sim\mathcal{N}(0,I)$, the policy $\bm{\pi}(s,c;\theta)$ can be modeled by a Gaussian Process with kernel $k_{\theta}((s,c),(s^{\prime},c^{\prime}))=\langle\phi_{\theta}(s,c),\phi_{\theta}(s^{\prime},c^{\prime})\rangle$.
 When coupled with this approximation, the conditional entropy objective in Equation 2 can be reformulated:
-$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}k_{\theta}((s_{t},c),X)[k_{\theta}(X,X)+\sigma_{\epsilon}^{2}I]^{-1}k_{\theta}(X,(s_{t},c)),$ (4)
-where $X$ is vector of states and tasks in $(c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$.
+$c_{n}=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\\ \tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\end{subarray}}\sum_{t=0}^{H-1}K(s_{t},c,X),$ (4)
+where $K(s_{t},\!c,\!X)\!:=\!k_{\theta}(x_{t},\!X)[k_{\theta}(X,\!X)\!+\!\sigma_{\epsilon}^{2}I]^{-1}k_{\theta}(X,\!x_{t})$, $x_{t}=(s_{t},c)$, and $X$ is the vector of states and tasks in $(c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$.
 As the collected dataset grows, the conditioning on previous trajectories $\tau_{1:n-1}$ can instead be addressed by fine-tuning the network’s parameters $\theta$ (e.g., through conventional gradient descent), resulting in updates in the embedding function $\phi_{\theta}$.
-Batch selection
-Standard practice in deep active learning prescribes the collection of batches of samples at each round (Gal et al., 2017; Sener & Savarese, 2018; Ducoffe & Precioso, 2018).
-This necessity is partially addressed by the fact that, in our setting, a single demonstration contains several samples for training the policy (as many as the length $H$ of the demonstration).
-Nevertheless, we can further leverage the GP approximation to select multiple demonstrations at each round.
-In fact, a simple recursive greedy selection can provide a constant factor approximation on the information gain objective (Krause & Golovin, 2014; Hübotter et al., 2024).
-In practice, when selecting $M$ demonstrations at the $n$-th round, we can select the $m$-th demonstration out of $M$ by applying the criterion in Equation 1, while making sure to condition entropy estimates on both the $m-1$ demonstrations already selected in this round, and the $M\cdot(n-1)$ demonstrations previously collected.
-Dealing with forgetting
-A fundamental issue with neural function approximation under shifts to the training distribution is known as forgetting (McCloskey & Cohen, 1989; French, 1999), as any further optimization may catastrophically perturb pre-trained parameters.
-Common strategies for its mitigation often involve rehearsal (Atkinson et al., 2021; Verwimp et al., 2021) or regularization (Kirkpatrick et al., 2017).
-Unfortunately, the former is not feasible in this setting due to lack of access to pre-training data, and the latter was not found to be empirically effective (see Appendix G).
-Scale and a diverse pre-training dataset can also mitigate forgetting (Dyer et al., 2022), but neither can be controlled during fine-tuning.
-Thus, we opt for a parameter isolating solution (Rusu et al., 2016; Yu et al., 2020), in which the task-space is partitioned (e.g.
-uniformly), and a copy of the fine-tuning parameters is stored and trained for each partition, thus avoiding negative interference.
-Inference can then be performed by selecting the parameter set through a nearest-neighbor lookup in task-space.
-This solution can be easily applied for limited task sets, and can be scaled to large task spaces through discretization schemes.
-While parameter isolation prevents constructive interference across tasks, we found it to bring a net benefit during fine-tuning.
-By combining the approximations required by AMF-GP with three additional design choices (namely kernel approximation, batch selection and parameter isolation), we obtain a method for active multi-task fine-tuning of policies parameterized via neural networks.
-We refer to this practical instantiation as AMF-NN.
-Figure 2: 2D integrator. Starting from the origin, each task involves reaching a given point on a circle, as shown in differently colored trajectories.
-5 Experiments
-The experiment section is designed to evaluate active multi-task fine-tuning and provide an empirical answer to several questions. We thus reserve a section to each of them.
-5.1 When is AMF beneficial?
-Uniform pre-training (12/12 tasks)
+No mismatch (12/12 tasks demonstrated)
   AMF-GP    Uniform
-Skewed pre-training (6/12 tasks)
+Mismatch (6/12 tasks demonstrated)
 Figure 2: Experiments in GP settings for a 2D integrator (see Figure 3).
 AMF-GP selects tasks that minimize the policy’s posterior entropy and improves the agent’s returns faster than uniform task sampling.
-In the middle, the improvement in final return over the baseline is greater when the pre-training distribution is skewed and includes fewer tasks.
-We report return and entropy curves for non-skewed and skewed pre-training (left and right, respectively).
-We report means and 90% simple bootstrap confidence intervals over 10 random seeds ; dots and crosses mark corresponding measurements.
+In the middle, the improvement in final return over the baseline is greater when the pre-training distribution does not perfectly match the evaluation distribution $\mu_{\mathcal{C}}$ and only demonstrates fewer tasks.
+We report return and entropy curves for non-mismatched and mismatched pre-training (left and right, respectively).
+We plot means and 90% simple bootstrap confidence intervals over 10 random seeds ; dots and crosses mark corresponding measurements.
+Dealing with forgetting
+Catastrophic forgetting is an inherent challenge to neural function approximation under shifts to the training distribution, as optimization over new data is non-local, and may undo learning progress (McCloskey & Cohen, 1989; French, 1999).
+This issue is critical in our setting, as actively guiding the fine-tuning distribution will necessarily accentuate the distribution shift from pre-training.
+Common strategies for its mitigation often involve rehearsal (Atkinson et al., 2021; Verwimp et al., 2021) or regularization (Kirkpatrick et al., 2017).
+Unfortunately, the former is not possible in this setting due to lack of access to pre-training data, and the latter was not found to be empirically effective (see Appendix H).
+Scale and a diverse pre-training dataset can also mitigate forgetting (Ramasesh et al., 2022), but neither can be controlled during fine-tuning.
+We thus propose a practical algorithmic solution to alleviate catastrophic forgetting when fine-tuning multi-task policies.
+Intuitively, we would like the policy to retain the skills it mastered during pre-training, while focusing on novel tasks during fine-tuning.
+Inspired by previous works in behavioral priors (Bagatella et al., 2022) and offline RL (Kumar et al., 2020), we propose retain a copy of the pre-trained policy, which we refer to as prior ($\bm{\pi}_{p}$).
+For a given state $s\in\mathcal{S}$ and task $c\in\mathcal{C}$, we can then linearly combine actions sampled from the fine-tuned policy $\bm{\pi}$ with those sampled from the prior: $a=\alpha(c)\hat{a}+(1-\alpha(c))\bar{a}$, where $\hat{a}\sim\bm{\pi}(\cdot|s,c),\bar{a}\sim\bm{\pi}_{\text{p}}(\cdot|s,c)$.
+Crucially, $\alpha(c)\in[0,1]$ is a task-dependent weight trained through gradient descend on an proxy behavior cloning loss, with an additional conservative penalty that encourages closeness to 00.
+As a result, actions will drift towards the fine-tuned policy’s output as soon as it robustly outperforms pre-training performance on a given task.
+Tasks which are not improved during fine-tuning will rely instead on samples from the prior, and will not be forgotten.
+We refer to this technique as Adaptive Prior, and present a detailed description and a comparison to continual learning techniques in Appendix H.
+By combining the approximations required by AMF-GP with the described solutions for estimating entropy and preventing catastrophic forgetting, we obtain a method for active multi-task fine-tuning of policies parameterized via neural networks, which we refer to as AMF-NN.
+5 Experiments
+The experiment section is designed to evaluate active multi-task fine-tuning and provide an empirical answer to several questions.
+We thus reserve a section to each of them.
+Additional evaluations are reported in Appendix D, E and F.
+5.1 When is AMF beneficial?
 When none of the assumptions listed in Section 4.1 is violated, AMF is guaranteed to converge to the optimal policy.
@@
 The agent is a pointmass initialized in the origin, and can directly control its 2D velocity, which is integrated over the past trajectory to return the current state.
-We can define an infinite task space, in which each task consists of reaching a point on a circle centered on the origin, and the agent is rewarded with the negative Euclidean distance to it.
+We can define a continuous task space, in which each task consists of reaching a point on a circle centered on the origin, and the agent is rewarded with the negative Euclidean distance to it.
 The evaluation distribution $\mu_{c}$ assigns equal probability to 12 points in different directions.
@@
 We model the policy as a Gaussian Process with a RBF kernel, and we condition it on a pre-training dataset of $12$ noisy demonstrations.
-We then collect $50$ additional demonstrations by running both AMF-GP and uniform sampling.
-We first consider a perfectly uniform pre-training regime, in which each evaluation task is demonstrated exactly once (Figure 3, left).
-As the policy’s entropy is minimized, AMF-GP increases the policy’s returns at a higher rate compared to uniform sampling of demonstrations.
-We then extend this evaluation to several pre-training distributions (Figure 3, middle), and compare the final performance of the two methods as the pre-training budget is allocated to a decreasing number of tasks.
-As the pre-training distribution becomes more skewed (e.g., when only 6/12 tasks are demonstrated in Figure 3, left), we observe that the performance gap between uniform task sampling and AMF-GP grows larger.
+We then collect $40$ additional demonstrations by running both AMF-GP and uniform sampling.
+Figure 3: 2D integrator. Starting from the origin, each task involves reaching a given point on a circle, as shown by differently colored trajectories.
+State-based
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
+01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
+01020Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
+01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
+Visual
+0510Rounds0.10.20.30.40.5Multi-task success rateKitchen, mismatch
+0510Rounds0.200.250.300.350.400.450.50Multi-task success rateKitchen, no mismatch
+0510Rounds0.30.40.50.60.7Multi-task success rateMetaworld, mismatch
+0510Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
+  AMF-NN   Uniform, with adaptive prior   Uniform
+Figure 4: AMF with neural policies in Frankakitchen (top) and Metaworld (bottom).
+Experiments are repeated for state and RGB inputs (left and right).
+We evaluate both mismatched and non-mismatched settings.
+AMF-NN is overall desirable, and highly beneficial for mismatched pre-training distributions.
+We report means and 90% simple bootstrap confidence intervals over 10 seeds.
+As a sanity check, we first consider a perfectly uniform pre-training regime, in which each evaluation task is demonstrated exactly once.
+The pre-training task distribution $\mu_{\mathcal{D}}\in\Delta(\mathcal{C})$ thus perfectly matches the evaluation task distribution $\mu_{\mathcal{C}}$ (Figure 2, left).
+As it actively minimizes the policy’s entropy, AMF-GP increases the policy’s returns slightly faster when compared to uniform sampling of demonstrations.
+We then extend this evaluation to more realistic pre-training distributions, characterised by a mismatch with respect to the evaluation distribution: $\mu_{\mathcal{C}}\neq\mu_{\mathcal{D}}$ (Figure 2, middle), and compare the final performance of the two methods as the pre-training budget is allocated to a decreasing number of tasks.
+As the pre-training distribution diverges from $\mu_{\mathcal{C}}$ (e.g., when only 6/12 tasks are demonstrated in Figure 2, right), we observe that the performance gap between uniform task sampling and AMF-GP grows larger.
 This is to be expected, as in this case the information gain from the next demonstration heavily depends on the queried task, and taking the $\operatorname*{arg\,max}$ of the criterion in Equation 1 is significantly better than choosing a random task.
 Intuitively, in this case, uniform sampling of tasks fails to reliably provide demonstrations for tasks that were observed less often during pre-training.
-State-based
-0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
-Visual
-0510Rounds0.200.250.300.350.40Multi-task success rateKitchen, uniform
-0510Rounds0.150.200.250.300.350.40Multi-task success rateKitchen, skewed
-0510Rounds0.50.60.70.80.9Multi-task success rateMetaworld, uniform
-0510Rounds0.40.60.8Multi-task success rateMetaworld, skewed
-  AMF-NN ($\sigma^{2}=10^{-4}$)   AMF-NN ($\sigma^{2}=10^{-3}$)    AMF-NN ($\sigma^{2}=10^{-2}$)
-  Uniform    Uniform, no parameter isolation
-Figure 4: AMF with neural policies in Frankakitchen (top) and Metaworld (bottom).
-Experiments are repeated for state and RGB inputs (left and right).
-We evaluate both uniform and skewed pre-training distribution.
-AMF-NN is overall desirable, and highly beneficial for skewed pre-training distributions.
-We report means and 90% simple bootstrap confidence intervals over 10 seeds.
 5.2 Can AMF scale to high-dimensional tasks?
@@
 In both environments, we evaluate AMF-NN when learning from state measurements, as well as from raw pixels.
-In the first case, the policy is simply parameterized through a MLP, while in the second the MLP receives the embedding of a pre-trained visual encoder (Nair et al., 2022b).
-The policy is pre-trained on $\approx 15$ total demonstrations, which we allocate either uniformly across all tasks, or only on half of them, reproducing the uniform and skewed regimes from the previous experiments.
-Afterwards, we apply AMF-NN for $10$ iterations, collecting $2$ demonstrations at each iteration.
+In the first case, the policy is simply parameterized through a MLP, while in the second the MLP receives the embedding of a pre-trained visual encoder (Nair et al., 2022).
+The policy is pre-trained on $\approx 15$ total demonstrations, which we allocate either uniformly across all tasks, or only on half of them, reproducing the non-mismatched and mismatched regimes from the previous experiments.
+Afterwards, when learning from state measurements, we apply AMF-NN for $20$ iterations, collecting one demonstrations at each iteration.
+To compensate for the increased complexity, in visual settings we instead provide 2 demonstrations per each task selected and only evaluate 10 iterations.
+01020Rounds0.30.40.50.60.70.80.9Multi-task success rateKitchen, mismatch
+01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
+  Last-layer gradient     Last-layer     Dropout
+Figure 5: AMF performance with alternative uncertainty estimation schemes.
+05Rounds0.30.40.50.60.7Multi-task success rateRobomimic, mismatch
+05Rounds0.30.40.50.60.7Multi-task success rateRobomimic, no mismatch
+  AMF-NN    Uniform, with adaptive prior    Uniform
+Figure 6: Evaluation on Robomimic. AMF-NN is widely applicable (e.g., to diffusion policies).
 Figure 4 reports average multi-task success rates at each iteration compared to a random uniform task selection scheme.
-For the baseline, we additionally report performance without parameter isolation, highlighting performance degradation due to negative gradient interference across tasks.
-We observe that AMF’s performance is on par or better than a uniform task sampling scheme when the prior was trained uniformly on all tasks.
-However, as reported in the previous section, AMF is very beneficial when the pre-training dataset does not uniformly cover the evaluation tasks.
-These trends are consistent across both environments, and both modalities.
-Finally, Figure 4 also highlights the effect of the only hyperparameter introduced by AMF-NN, namely the noise parameter $\sigma^{2}$ in the GP approximation.
-While selecting an excessively high noise level results in slightly greedier behavior and premature convergence, we overall observe that the method is not particularly sensitive to this hyperparameter, performing fairly reliably across 2 orders of magnitude.
-5.3 Is batch-wise task selection important for AMF?
-In the GP setting, policy fine-tuning and conditioning for batch-wise selection correspond.
-Thus, it is not necessary to collect batches of demonstrations.
-This is not the case for AMF-NN.
-On one hand, collecting batches of training data is computationally beneficial since training can be parallelized effectively (Sener & Savarese, 2018).
-On the other hand, each query already returns several samples in our setting, which could make batch selection unnecessary.
-We thus set out to empirically validate which batch size is desirable for active data collection in our setting.
-4.04.55.05.56.06.5Area under success rate curveKitchen, uniform
-4.04.55.05.56.06.5Area under success rate curveKitchen, skewed
-$B\!=1$ $B\!=2$ $B\!=4$ Uniform
-Uniform, no parameter isolation
-Figure 5: Influence of batch size $B$ over area under success rate curve with 20 demonstrations.
-0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
-0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
-  Last-layer gradient
-  Last-layer    Dropout
-Figure 6: AMF performance with alternative uncertainty estimation schemes.
-While keeping the total budget fixed to 20 demonstrations, we evaluate AMF-NN with batch sizes $B$ spanning from 1 to 4, and report the area under the success rate curve in Figure 5.
-For convenience, we report the same metric for the uniform selection baselines from Figure 4, for which the data selection strategy does not depend on the batch size.
-We find that larger batch sizes are not necessarily desirable.
+For the baseline, we additionally report performance with an adaptive prior, highlighting how this solution prevents performance degradation due to catastrophic forgetting of pre-training demonstrations.
+As reported in the previous section, AMF is beneficial under distribution mismatch, in which the pre-training dataset does not exactly cover the evaluation distribution $\mu_{\mathcal{C}}$.
+As a sanity check, we also observe that, when pre-training demonstrates all tasks equally and a uniform task allocation would be very effective, AMF’s performance matches this naive baseline.
+These trends are consistent across both environments, and both modalities. For a qualitative analysis of the strategy induced by AMF, we refer to Appendix I and J.
 5.3 How do uncertainty estimates for AMF compare?
 As entropy estimation is at the core of AMF-NN, we additionally compare the adopted GP approximation with loss-gradient embeddings to other approaches from the literature.
-In praticular, we also consider an alternative GP approximation using last-layer embeddings (Holzmüller et al., 2023), as well as test-time Dropout (Loquercio et al., 2020).
-For the latter, each batch is simply filled with demonstrations from the task maximizing prior entropy, that is $\operatorname*{arg\,max}_{c\in\mathcal{C}}\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;(s_{0},\dots)\sim\bm{\tau}(c)}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid\tau_{1:n-1})$.
+In particular, we also consider an alternative GP approximation using last-layer embeddings (Holzmüller et al., 2023), as well as test-time Dropout (Loquercio et al., 2020).
+The latter simply selects the task maximizing prior entropy, that is $\operatorname*{arg\,max}_{c\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid\tau_{1:n-1})$, with $\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})$ and $(s_{0},\dots)\sim\bm{\tau}(c)$.
 Both of these schemes are in practice desirable, as they do not require access to action labels.
-However, we observe that these two schemes are prone to early convergence to suboptimal task choices, or are less effective in driving task selection.
-Hence, as shown in Figure 6, multi-task performance is in general lower, suggesting that the entropy estimation technique is crucial to AMF-NN.
+However, we observe that these two schemes are less effective in driving task selection.
+Hence, as shown in Figure 6 (and in Appendix G), multi-task performance is in general lower, suggesting that the entropy estimation technique is important for AMF-NN.
 5.4 Can AMF be applied to off-the-shelf models?
-  AMF-NN    Uniform
-  Uniform, no parameter isolation
-024Rounds0.400.450.500.550.600.65Multi-task success rateWidowX
-Figure 7: Evaluation on life-like WidowX tasks. AMF-NN can be applied to large-scale settings.
 As AMF-NN has minimal requirements (essentially, access to a differentiable pre-trained prior is sufficient), it should be widely applicable.
-In this section, we investigate scaling our evaluation to recently published open-source generalist policies.
-For this purpose, we chose Octo (Octo Model Team et al., 2024).
-This model relies on a transformer backbone for integrating multimodal information (in the form of state sensors, camera images and text or RGB task descriptions), and uses a diffusion-based policy head for action prediction (Chi et al., 2023).
-For computational reasons, we will focus on fine-tuning the action head alone.
-Octo is pre-trained on a large-scale real-world robotic dataset (Collaboration, 2023), and is thus designed for inference on physical hardware.
-Nonetheless, a recently proposed evaluation suite enables simulated evaluations that statistically correlate with real-world results (Li et al., 2024).
-We thus collect rollouts from a pre-trained Octo on the WidowX tasks, and filter them to only include successes, akin to self-distillation schemes Bousmalis et al.
-(2024).
-On availability of such self-supervised demonstrations, we then apply AMF-NN for $4$ iterations, providing $4$ demonstrations in each round.
-The results are reported in Figure 7.
-As all evaluation tasks are largely demonstrated in the pre-training dataset (Collaboration, 2023), we find that AMF-NN does not improve significantly upon uniform task collection, confirming the trend we observed for uniform pre-training distributions in Figure 4.
-Nonetheless, we observe that it constitutes an effective method for data selection, and can be applied as a drop-in replacement for fine-tuning of off-the-shelf models.
-[^3] This evaluation also reports an interesting trend, that is a vast reduction in catastrophic forgetting, to the point that parameter isolation is not necessary.
-This anecdotal evidence can be seen as an instance of a general trend of mitigated catastrophic forgetting in large models (Dyer et al., 2022).
+In this section we investigate further scaling our evaluation to more complex tasks, multi-modal demonstrators and modern policy classes.
+We consider the Robomimic benchmark, which involves four long-horizon, precise manipulation tasks (up to $\approx$700 steps).
+While experiments in previous sections rely on demonstrations collected by scripted policies or RL agents, expert trajectories are in this case provided by humans; due to the increased scale, we sample 20 demonstrations for the task selected at each iteration.
+Finally, we fine-tune a larger generative model, namely a diffusion policy (Chi et al., 2023), which remains compatible with AMF-NN.
+Despite the change in data source and architectures, the results we observe in Figure 6 are consistent with those reported in previous settings: active fine-tuning and an adaptive prior are overall helpful, especially when the pre-training distribution does not match the evaluation distribution $\mu_{\mathcal{C}}$.
 6 Discussion
@@
 Second, we found the performance of AMF to depend naturally on the pre-training data distribution.
-While AMF induces efficient learning for skewed pre-training distributions, it naturally brings more modest gains when the pre-trained policy is equally capable for all tasks, and uniform task sampling is sufficient.
+While AMF induces efficient learning for mismatched pre-training distributions, it naturally brings more modest gains when the pre-trained policy is equally capable for all tasks, and uniform task sampling is sufficient.
 On top of addressing the current limitations, this work suggests multiple interesting directions.
-An extensive empirical evaluation of active fine-tuning with large-scale generalist policies is clearly desirable, but remains infeasible at the moment due to the scarce availability of open-source models and benchmarks.
+An extensive empirical evaluation of active fine-tuning with large-scale generalist policies is clearly desirable, but remains infeasible at the moment due to the scarce availability of open-source calibrated benchmarks.
 Another future research direction would involve direct estimation of the RL objective, thus removing the dependence on non-equivalent BC proxy objectives.
@@
 815943, and from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545.
-Reproducibility Statement
-Appendix I describes implementation details and hyperparameters across all experiments. We additionally open-source a clean and commented implementation of AMF-GP and AMF-NN at the anonymous website.
+Impact Statement
+This paper formally introduces a framework for active data selection in MDPs, and investigates its applicability in diverse settings.
+Although this direction as a whole might eventually have an effect on labor and automation, we do not foresee a direct impact of our work in terms of societal consequences.
 Appendix A Performance guarantees under regularity assumptions
@@
 We thus introduce the following:
-Assumption 1.
+Assumption A.1.
 (Regular, noisy policy) We assume that the optimal policy $\pi^{\star}\sim GP(\mu,k)$ with known mean function $\mu$ and kernel $k$.
@@
 For this reason, we require the following two sets of assumptions to ensure that information gained along empirical trajectories is not arbitrarily smaller than the expected one.
-Assumption 2.
+Assumption A.2.
 (Lipschitz, bounded MDP and policy) Given the contextual MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{C},P,R,\gamma,\mu_{0})$ and the noisy expert $\bm{\tilde{\pi}}$ we assume that, for every $\{(s,c,a),(s^{\prime},c^{\prime},a^{\prime})\}\subseteq\mathcal{S\times C\times A}$:
@@
 We note that smoothness of the noisy expert is guaranteed by construction if the expert $\pi^{\star}$ is $L_{\pi}$-smooth.
-Assumption 3.
+Assumption A.3.
 (Smooth MI) For every pair of sequences of trajectories $\{\tau_{1:n-1},\tau^{\prime}_{1:n-1}\}\subseteq(\mathcal{S\times A})^{H(n-1)}$, $(s,c)\in\mathcal{S\times C}$, $c_{1:n-1}\in\mathcal{C}^{n-1}$, $\tilde{\tau}\in(\mathcal{S\times A})^{H}$ and $c_{n}\in\mathcal{C}$, we assume that the mutual information at step $n$ is $L_{I}$-smooth with respect to the mean square deviation of collected trajectories:
@@
 A.2 Proof
-We first prove that, under Assumptions 2 and 3, the effect of stochasticity on the mutual information at step $n$ is bounded.
-Lemma 1.
-Let Assumptions 2 and 3 hold.
+We first prove that, under Assumptions A.2 and A.3, the effect of stochasticity on the mutual information at step $n$ is bounded.
+Lemma A.4.
+Let Assumptions A.2 and A.3 hold.
 Fix a sequence of tasks $c_{1:n-1}$ and consider two arbitrary sequences of trajectories $\tau_{1:n-1}$ and $\tau^{\prime}_{1:n-1}$ sampled from $\bm{\tau}(c_{1:n-1})$.
@@
 Proof.
-Under Assumption 3 it is sufficient to show that stochasticity in the MDP does not cause the demonstrator’s trajectories to deviate excessively.
-This is a direct consequence of smoothness and boundedness, which we assume in Assumption 2, and can be shown by induction.
+Under Assumption A.3 it is sufficient to show that stochasticity in the MDP does not cause the demonstrator’s trajectories to deviate excessively.
+This is a direct consequence of smoothness and boundedness, which we assume in Assumption A.2, and can be shown by induction.
 Let us fix a task $c_{n}\in\mathcal{C}$ and consider two trajectories $\tau,\tau^{\prime}\sim\bm{\tau}(c_{n})$.
@@
 $:=A\epsilon_{t}+B,$ (16)
-where Lemma 9 was used in (i) and (ii); Assumption 2 and the fact that $c_{t}=c^{\prime}_{t}$ were used through the rest of the derivation. The recurrence relation can be easily unrolled as
+where Lemma N.2 was used in (i) and (ii); Assumption A.2 and the fact that $c_{t}=c^{\prime}_{t}$ were used through the rest of the derivation. The recurrence relation can be easily unrolled as
 $\epsilon_{t}$ $\leq A^{t}\epsilon_{0}+\sum_{i=0}^{t-1}A^{i}B$ (17)
@@
 thus bounding the L2 distances between states at each step of the trajectory $\epsilon_{t}=\|s_{t}-s^{\prime}_{t}\|_{2}$.
-We note that the distance between actions can also be easily bound by Lemma 9: $\|a_{t}-a^{\prime}_{t}\|_{2}\leq L_{\pi}\epsilon_{t}+2\epsilon_{\pi}$.
+We note that the distance between actions can also be easily bound by Lemma N.2: $\|a_{t}-a^{\prime}_{t}\|_{2}\leq L_{\pi}\epsilon_{t}+2\epsilon_{\pi}$.
 This can in turn be related to distances over trajectories.
@@
 $\leq 8H^{\frac{3}{2}}(1+\max(L_{P},L_{\pi}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P}).$ (35)
-Having obtained an upper bound on the distance between sequences of trajectories, the result follows naturally from smoothness of mutual information according to Assumption 3.
+Having obtained an upper bound on the distance between sequences of trajectories, the result follows naturally from smoothness of mutual information according to Assumption A.3.
 We can now focus on the main result. We start by introducing an important measure, quantifying the maximum information gain at each round:
@@
 As common in the literature (Bogunovic et al., 2016; Kothawade et al., 2020; Hübotter et al., 2024), we make a standard assumption on diminishing informativeness.
-Assumption 4.
+Assumption A.5.
 For each $n,i\in\mathbb{N}$ with $i\leq n$, the maximum information gain at round $n$ is not greater than the maximum information gain at round $i$:
@@
 This can be leveraged to show that the expected mutual information is sublinear in the number of rounds $n$.
-From this point, we overload the notation and allow policies (e.g.
-$\bm{\pi}$) to map vector to random vectors, that is $\bm{\pi}((x_{0},\dots,x_{n-1}))=(\bm{\pi}(x_{0}),\dots,\bm{\pi}(x_{n-1}))$ for $(x_{0},\dots,x_{n-1})\in(\mathcal{S\times C})^{n}$.
-Lemma 2.
-Under Assumptions 1 and 4, if $(c_{0},\dots,c_{n})$ follows the criterion in Equation 6, then $\Gamma_{n}\leq\frac{H}{n}\gamma_{(Hn)}$, where $\gamma_{(Hn)}=\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|\leq Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$.
+From this point, we overload the notation and allow policies (e.g., $\bm{\pi}$) to map vector to random vectors, that is $\bm{\pi}((x_{0},\dots,x_{n-1}))=(\bm{\pi}(x_{0}),\dots,\bm{\pi}(x_{n-1}))$ for $(x_{0},\dots,x_{n-1})\in(\mathcal{S\times C})^{n}$.
+Lemma A.6.
+Under Assumptions A.1 and A.5, if $(c_{0},\dots,c_{n})$ follows the criterion in Equation 6, then $\Gamma_{n}\leq\frac{H}{n}\gamma_{(Hn)}$, where $\gamma_{(Hn)}=\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|\leq Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$.
 Proof.
@@
 $=\frac{H}{n}\gamma_{(Hn)}$ (46)
-where (i) follows from Assumption 4, (ii) follows from Equation 6, (iii) is due to the chain rule of mutual information.
+where (i) follows from Assumption A.5, (ii) follows from Equation 6, (iii) is due to the chain rule of mutual information.
 We note that $\gamma_{n}=\max_{X\subseteq\mathcal{S}\times\mathcal{C},\;|X|\leq n}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ is sublinear for a large class of GPs.
@@
 This bound on expected round-wise mutual information can then be leveraged to describe how the total variance shrinks over rounds.
-Lemma 3.
-(Uniform convergence of marginal variance, following Hübotter et al. (2024)) Under Assumption 1, 2 and 3, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
+Lemma A.7.
+(Uniform convergence of marginal variance, following Hübotter et al. (2024)) Under Assumption A.1, A.2 and A.3, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
 $\sigma_{n}^{2}(s,c)\leq(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau_{\text{min}}^{2}},$
@@
 $=(1+\epsilon_{n})\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau_{\text{min}}^{2}}.$ (64)
-where (i) follows from Lemma 8 and monotonicity of variance, (ii) holds as the state $s$ is within the support of $\bm{\tau}(\cdot|c)$, and (iii) follows from Lemma 1 as the difference between the expected mutual information and the mutual information for a realized trajectory is less than the difference in mutual information for two arbitrary realized trajectories.
+where (i) follows from Lemma N.1 and monotonicity of variance, (ii) holds as the state $s$ is within the support of $\bm{\tau}(\cdot|c)$, and (iii) follows from Lemma A.4 as the difference between the expected mutual information and the mutual information for a realized trajectory is less than the difference in mutual information for two arbitrary realized trajectories.
 ∎
@@
 Without loss of generality we will assume that the prior variance is bounded by $\mathrm{Var}[\bm{\pi}(s,c)]\leq 1$.
-Lemma 4.
+Lemma A.8.
 (Well-calibrated confidence intervals, following Abbasi-Yadkori (2013)) Pick $\delta\in(0,1)$.
@@
 We note that $\beta_{n}(\delta)$ depends on $\gamma_{(Hn)}$ as $Hn$ samples from the demonstrator’s policy are collected up to round $n$.
-Combining Lemmas 3 and 4 we easily get for all $(s,c)\in\mathcal{S\times C}$ and $n\geq 0$ with probability $1-\delta$:
+Combining Lemmas A.7 and A.8 we easily get for all $(s,c)\in\mathcal{S\times C}$ and $n\geq 0$ with probability $1-\delta$:
 […]
@@
 If smoothness holds, it is easy to connect the divergences in action space to performance gaps (Maran et al., 2023).
-Lemma 5.
+Lemma A.9.
 Let $\pi$ and $\pi^{\prime}$ denote two deterministic policies. If the state-action value function $Q^{\pi^{\prime}}$ is $L_{Q^{\pi^{\prime}}}$-Lipschitz continuous, then:
@@
 So far, we have shown rates of convergence for the imitator, and connected its error to performance. Our main formal result can be shown by coordinating the lemmas so far presented.
-Theorem 3.
-(Performance guarantees for active multi-task BC) Let Assumptions 2, 3 and 4 hold.
+Theorem A.10.
+(Performance guarantees for active multi-task BC) Let Assumptions A.2, A.3 and A.5 hold.
 Pick $\delta\in(0,1)$.
@@
 If each demonstrated task is selected according to the criterion in Equation 1, then with probability at least $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
-$J^{\pi^{\star}}-J^{\pi_{n}}\leq\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}=O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
+$J^{\pi^{\star}}-J^{\pi_{n}}\leq\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}=O(\gamma_{(Hn)})/\sqrt{n},$
 where $\epsilon_{n}=8H^{\frac{3}{2}}(1+\max(L_{\pi},L_{P}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$.
@@
 where (i) is due to the fact that $J^{\pi^{\star}}\geq J^{\pi}$ for any policy $\pi$, and the expectation fades due to uniform convergence.
-The only terms with a dependency on $n$ are $\beta^{\prime}_{n}(\delta)=O(\gamma_{(Hn)}^{\frac{1}{2}})$ and $\Gamma_{n}=O(\gamma_{(Hn)}n^{-1})$, which can be combined in the asymptotic notation in the Theorem.
-If $\gamma_{n}=O(\log n)$, then $J^{\pi^{\star}}-J^{\pi}=O(n^{-\frac{1}{2}}\log n)\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$.
+The only terms with a dependency on $n$ are $\beta^{\prime}_{n}(\delta)=O(\gamma_{(Hn)}^{\frac{1}{2}})$ and $\Gamma_{n}=O(\gamma_{(Hn)})/n$, which can be combined in the asymptotic notation in the Theorem.
+If $\gamma_{n}=O(\log n)$, then $J^{\pi^{\star}}-J^{\pi}=O(\log n)/\sqrt{n}\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$.
 For a summary of magnitudes of $\gamma_{n}$ for common kernels, we refer to Table 3 in Hübotter et al.
@@
 Appendix B Guarantees in non-Lipschitz MDPs
-The main result reported in Theorem 3 provides anytime guarantees on the agent’s performance, assuming smoothness in the MDP.
+The main result reported in Theorem A.10 provides anytime guarantees on the agent’s performance, assuming smoothness in the MDP.
 However, it is possible to replace this assumption with a weaker one, at the cost of only retaining guarantees in expectation.
@@
 (2023).
-Assumption 5.
+Assumption B.1.
 The noise distribution $\bm{\epsilon}$ is $L_{\ell}$-TV-Lipschitz continuous.
 This assumption is satisfied by a large class of Gaussian and sub-Gaussian distributions (Maran et al., 2023).
-We can build upon Assumption 4 and Lemma 2, and start by providing a weaker version of Lemma 3.
-Lemma 6.
-(Uniform convergence of marginal variance in expectation) Under Assumption 1, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
+We can build upon Assumption A.5 and Lemma A.6, and start by providing a weaker version of Lemma A.7.
+Lemma B.2.
+(Uniform convergence of marginal variance in expectation) Under Assumption A.1, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
 $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\sigma^{2}_{n}(s,c)\leq\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau^{2}_{\text{min}}},$
@@
 Proof.
-We resume from Inequality 60 in the proof of Lemma 3:
+We resume from Inequality 60 in the proof of Lemma A.7:
 $\sigma_{n}^{2}(s,c)\leq\frac{2\bar{\sigma}^{2}}{\tau^{2}_{\text{min}}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c}\\ \tau\sim\bm{\tau}(c)\\ (s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}\mathcal{I}(\bm{\pi}(s_{t},c);\bm{\tilde{\pi}}(\tau,c)\mid c_{1:n},\tau_{1:n})$ (79)
@@
 $=\frac{2\bar{\sigma}^{2}\Gamma_{n}}{\tau^{2}_{\text{min}}}.$ (82)
-Having bounded variance at each round, this time in expectation, we can invoke Lemma 4 to bound the expected distance to the optimal policy with high probability:
+Having bounded variance at each round, this time in expectation, we can invoke Lemma A.8 to bound the expected distance to the optimal policy with high probability:
 $\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})}\|\pi^{\star}(s,c)-\mu_{n}(s,c)\|_{1}\leq\beta^{\prime}_{n}(\delta)\|\bar{\sigma}\|_{1}\Big{(}\frac{2\Gamma_{n}}{\tau^{2}_{\text{min}}}\Big{)}^{\frac{1}{2}}$ (83)
@@
 (2023), and refer to the original work for the proof.
-Lemma 7.
-Let $\pi^{\star}$, $\bm{\tilde{\pi}}$ and $\pi$ denote the expert, noisy expert and imitator policy, respectively. If Assumption 5 holds, then:
+Lemma B.3.
+Let $\pi^{\star}$, $\bm{\tilde{\pi}}$ and $\pi$ denote the expert, noisy expert and imitator policy, respectively. If Assumption B.1 holds, then:
 $J^{\bm{\tilde{\pi}}}-J^{\pi}\leq\frac{2L_{\ell}Q_{\text{max}}}{1-\gamma}\mathbb{E}_{s\sim\mu_{\bm{\tilde{\pi}}}}[\mathcal{W}(\pi^{\star}(\cdot\mid s),\pi(\cdot\mid s))],$
@@
 Results in Figure 4 are computed over two representative pre-training distributions: one allocating pre-training demonstrations uniformly over all tasks, the other one only demonstrating the first two tasks.
-We report these results again, and compare them with those for several other pre-training distributions.
-In particular, we evaluate a family of skewed priors which are only trained on one or two tasks.
+We report these results again, and compare them with those for several other pre-training distributions in which tasks have been shuffled.
 Results are reported for FrankaKitchen in Figure 9 and for Metaworld in Figure 9, and are consistent with patterns observed in Figure 4.
-0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
 $[8,8,0,0,0]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
 $[0,8,8,0,0]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
 $[0,0,8,8,0]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
+01020Rounds0.30.40.50.60.70.80.9Multi-task success rateKitchen, mismatch
 $[0,0,0,8,8]$
-0510Rounds0.20.40.60.8Multi-task success rateKitchen, skewed
+01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
 $[3,3,3,3,3]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[16,0,0,0,0]$
-  AMF-NN ($\sigma=1e^{-3}$)    Uniform    Uniform, no parameter isolation
+  AMF-NN    Uniform, with adaptive prior    Uniform
 Figure 8: Additional results for AMF-NN in FrankaKitchen with state inputs.
 We evaluate several allocations of the pre-training demonstrations, as labeled below each plot (e.g., the label $[8,8,0,0,0]$ indicates that 8 demonstrations were provided for each of the first two tasks each, and none for the remaining tasks).
-0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
+01020Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
 $[8,8,0,0]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
+01020Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
 $[0,8,8,0]$
-0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
+01020Rounds0.40.50.60.7Multi-task success rateMetaworld, mismatch
 $[0,0,8,8]$
-0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
+01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
 $[2,2,2,2]$
-0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
-$[16,0,0,0]$
-  AMF-NN ($\sigma=1e^{-3}$)   Uniform    Uniform, no parameter isolation
+  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    Uniform
 Figure 9: Additional results for AMF-NN in Metaworld with state inputs.
 We evaluate several allocations of the pre-training demonstrations, as labeled below each plot (e.g., the label $[8,8,0,0]$ indicates that 8 demonstrations were provided for each of the first two tasks each, and none for the remaining tasks).
-Moreover, we also report an ablation over the choice of batch size for Metaworld in Figure 11, thus complementing the one reported in Section 5.3.
-We observe a slight upward trend favoring larger batch size in the case of an uniform prior, but smaller batches remain overall desirable.
-34567Area under success rate curveMetaworld, uniform
-34567Area under success rate curveMetaworld, skewed
-$B\!=1$
-$B\!=2$
-$B\!=4$
+Appendix F AMF with generalist policies
+  AMF-NN
+  Uniform, with adaptive prior
   Uniform
-Uniform, no
-parameter isolation
-Figure 11: Influence of batch size $B$ over area under success rate curve with a budget of 20 demonstrations in Metaworld.
-Appendix F Uncertainty ablation
-Section 5.4 evaluates alternative uncertainty quantification schemes in FrankaKitchen for two pre-training distributions.
-This Section extends these results to include results for Metaworld, and for several other pre-training settings (see Figures 12 and 13).
+024Rounds40506070Multi-task success rateWidowX
+Figure 10: Evaluation on life-like WidowX tasks. AMF-NN can be applied to large-scale settings.
+This section investigates scaling our evaluation to recently published open-source generalist policies.
+For this purpose, we choose Octo (Octo Model Team et al., 2024).
+This model relies on a transformer backbone for integrating multimodal information (in the form of state sensors, camera images and text or RGB task descriptions), and uses a diffusion-based policy head for action prediction (Chi et al., 2023).
+For computational reasons, we will focus on fine-tuning the action head alone.
+Octo is pre-trained on a large-scale real-world robotic dataset (Collaboration, 2023), and is thus designed for inference on physical hardware.
+Nonetheless, a recently proposed evaluation suite enables simulated evaluations that statistically correlate with real-world results (Li et al., 2024).
+We thus collect rollouts from a pre-trained Octo agent on the WidowX tasks, and filter them to only include successes, akin to self-distillation schemes (Bousmalis et al., 2024).
+On availability of such self-supervised demonstrations, we then apply AMF-NN for $5$ iterations, providing $2$ demonstrations in each round.
+The results are reported in Figure 10.
+As all evaluation tasks are largely demonstrated in the pre-training dataset (Collaboration, 2023), and the evaluation is significantly noisier with respect to other benchmarks, we find that the performance of AMF-NN falls within confidence intervals of uniform task collection, confirming the trend we observed for uniform pre-training distributions in Figure 4.
+Nonetheless, we observe that it constitutes an effective method for data selection, and can be applied as a drop-in replacement for fine-tuning of off-the-shelf models.
+[^3] This evaluation also reports an interesting trend, that is a vast reduction in catastrophic forgetting, to the point that an adaptive prior is almost not necessary.
+This anecdotal evidence can be seen as an instance of a general trend of mitigated catastrophic forgetting in large models (Ramasesh et al., 2022).
+Appendix G Uncertainty ablation
+01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
+01020Rounds0.40.50.60.70.8Multi-task success rateMetaworld, mismatch
+  Last-layer gradient    Last-layer    Dropout
+Figure 11: Additional results for AMF-NN in Metaworld with state inputs and different uncertainty quantification techniques.
+Section 5.3 evaluates alternative uncertainty quantification schemes in FrankaKitchen for two representative pre-training distributions.
+This Section extends these results to include results for Metaworld (see Figure 11).
 Results are consistent with those so far reported, suggesting that loss gradient embeddings are an important component for the empirical performance of AMF-NN.
-0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
-$[3,3,3,3,3]$
-0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[8,8,0,0,0]$
-0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[0,8,8,0,0]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[0,0,8,8,0]$
-0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[0,0,0,8,8]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-$[16,0,0,0,0]$
-  Last-layer gradient    Last-layer    Dropout
-Figure 12: Additional results for AMF-NN in FrankaKitchen with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
-0510Rounds0.550.600.650.700.75Multi-task success rateMetaworld, uniform
-$[4,4,4,4]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
-$[8,8,0,0]$
-0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
-$[0,8,8,0]$
-0510Rounds0.50.60.70.8Multi-task success rateMetaworld, skewed
-$[0,0,8,8]$
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
-$[16,0,0,0]$
-  Last-layer gradient    Last-layer    Dropout
-Figure 13: Additional results for AMF-NN in Metaworld with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
-Appendix G Mitigating forgetting
+Appendix H Mitigating forgetting
 The ability of neural networks to adapt to shifts in training distribution while retaining information is an important object of interest in lifelong and continual learning (Wang et al., 2024).
@@
 Hence, the fine-tuning procedure should be careful not to disrupt this ability.
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
+01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
+01020Rounds0.20.40.60.8Multi-task success rateMetaworld, mismatch
+01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
+  Adaptive Prior    EWC    L2 regularization    Baseline
+Figure 12: Performance of AMF-NN with several techniques to mitigate forgetting. Darker shades represent stronger regularization coefficients. L2, EWC regularization are not effective in this setting.
 Several methods aimed at mitigating forgetting can be traced back to rehearsal (Riemer et al., 2019; Chaudhry et al., 2019) and regularization (Kirkpatrick et al., 2017) strategies.
@@
 The latter can be seen as a more nuanced version of the former, which adaptively scales the regularization strength according to the curvature of the loss landscape.
-Unfortunately, we found both methods to be insufficient in our setting, as reported in Figure 14.
-When coupled with large regularization weights, their asymptotical performance is significantly limited.
+Unfortunately, we find these methods to be insufficient in our setting, as reported in Figure 12.
+When coupled with large regularization weights, the asymptotical performance of L2-regularization and EWC is significantly limited.
 When regularization weights are too low, they recover the performance of a naive baseline.
 Intermediate values were found to interpolate between the two behaviors, without addressing the forgetting issue.
-While less scalable, we found hard parameter isolation to be the only effective solution amongst the one we tested.
-As mitigating forgetting in neural network is an orthogonal direction to the main topic of this work, we adopt this solution, and expect that further developments in continual learning will be applicable in our setting.
-0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, uniform
-0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
-0510Rounds0.40.50.60.7Multi-task success rateMetaworld, uniform
-0510Rounds0.20.40.6Multi-task success rateMetaworld, skewed
-  Parameter isolation    EWC    L2 regularization    Baseline
-Figure 14: Performance of uniform task sampling with several techniques to mitigate forgetting.
-Darker shades represent stronger regularization coefficients.
-L2 and EWC regularization are not effective in this setting.
-Appendix H Criterion vs returns
+As a result, the policy quickly loses information on its pre-training tasks, thus rendering adaptive data selection strategies ineffective.
+This motivates our adoption of a novel technique to alleviate forgetting, which we refer to as Adaptive Prior.
+Inspired by previous works in behavioral priors (Bagatella et al., 2022) and offline RL (Kumar et al., 2020), we linearly combine the policy’s output with that of a prior (in practice, a frozen copy $\bm{\pi}_{p}$ of the pre-trained policy).
+When an action needs to be sampled, we instead output a linear combination of actions sampled from the fine-tuned policy $\bm{\pi}$ and the prior $\bm{\pi}_{p}$: for a given state-task pair $(s,c)\in\mathcal{S\times C}$, the action selected is $a=\alpha(c)\hat{a}+(1-\alpha(c))\bar{a}$, where $\hat{a}\sim\bm{\pi}(\cdot|s,c)$ and $\bar{a}\sim\bm{\pi}_{p}(\cdot|s,c)$.
+The weight $\alpha(c)\in[0,1]$ is task-dependent and learnable: ideally, it would be high when the fine-tuned policy is more suitable for the task, and low when the prior is instead preferable.
+In practice, it may be a per-task learnable parameter for finite task spaces, or the output of a parametrized function (e.g., a neural network) in continuous task spaces.
+While this formulation departs from existing ones (Bagatella et al., 2022), in which the action is instead sampled from a mixture of policy and prior, it enables a simple update rule for $\alpha$, which remains applicable when the policy class does not allow straightforward likelihood evaluation (e.g., Diffusion policies).
+Under the assumption that both $\bm{\pi}$ and $\bm{\pi}_{p}$ are isotropic Gaussians, the BC loss computed with respect to the linear combinations of actions simplifies to a simple mean squared error:
+$\mathcal{L}^{\alpha}_{\text{BC}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\|a_{t}^{i}-(\alpha(c_{i})\hat{a}+(1-\alpha(c_{i}))\bar{a})\|_{2},$ (93)
+where $\hat{a}\sim\bm{\pi}(\cdot|s_{t}^{i},c_{i}),\bar{a}\sim\bm{\pi}_{p}(\cdot|s_{t}^{i},c_{i})$ and $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ is the dataset of $N$ task-conditioned, $H$-length trajectories with task labels $c_{1:N}$.
+As this loss function is differentiable with respect to $\alpha(c)$, the weight can be easily trained through gradient descent.
+Intuitively, this gradient pushes weights up for tasks in which the fine-tuned policy is more accurate than the pre-trained one.
+As $\bm{\pi}$ is fine-tuned on the dataset used for estimating $\mathcal{L}_{\text{BC}}^{\alpha}$, its performance can be easily overestimated.
+Thus, we propose to introduce a conservative penalty term, which encourages the weight to be updated only if the fine-tuned policy $\bm{\pi}$ significantly outperforms the prior $\bm{\pi}_{p}$: $\mathcal{L}^{\alpha}=\mathcal{L}^{\alpha}_{\text{BC}}+\beta\frac{1}{N}\sum_{i=1}^{N}\alpha(c_{i}).$ We remark that, in our empirical evaluation, the Gaussian assumption is often violated; nevertheless, we found this update rule to be empirically effective across models and tasks.
+To illustrate, we visualize trajectories of mixing weights $\alpha(c)$ during fine-tuning in Figure 13.
+We consider a mismatched setting, in which the pre-training distribution largely demonstrates the first two tasks in Kitchen, and does not cover the remaining tree.
+During early training, AMF-NN mostly samples demonstrations for the last three tasks.
+Thus, $\bm{\pi}$ quickly improves, and the respective mixing coefficients converge to 1.
+On the other hand, $\bm{\pi}_{p}$ outperforms $\bm{\pi}$ on the first two tasks for the majority of training, $\alpha$ remains low and evaluation performance does not drop as actions are largely sampled from $\bm{\pi}_{p}$.
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 13: Success rates (blue) and mixing coefficients $\alpha(c)$ (grey) over training in Kitchen with skewed pre-training.
+$\alpha$ grows quickly for novel tasks, and more conservatively for those that were demonstrated during pre-training.
+Appendix I Does AMF rebalance demonstration counts?
+In discrete task spaces, counting the number of demonstrations for each task is possible.
+In this case, a naive data selection strategy would simply request demonstrations for tasks that have been demonstrated the least in the past.
+If all tasks require a similar amount of demonstrations, this would empirically perform very well.
+In our setting, however, data selection algorithms do not have knowledge of pre-training data.
+For this reason, a count could only be kept with respect to the fine-tuning demonstrations: actively balancing this count would lead to a near-uniform task selection, and recover the performance of uniform sampling in expectation.
+01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
+01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
+01020Rounds0.20.40.60.8Multi-task success rateMetaworld, mismatch
+01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
+  AMF-NN    Rebalancing    Uniform, with adaptive prior    Uniform
+Figure 14: Extended results from Figure 4, including a privileged “rebalancing” baseline.
+Nevertheless, we implement this “rebalancing” criterion as a privileged baseline, which assumes access to the pre-training task distribution.
+We evaluate it in the standard settings for AMF-NN from Figure 4.
+In Figure 14, we observe that AMF-NN is able to match the performance of this privileged baseline, despite having no knowledge of the pretraining distribution.
+This implies that AMF can infer information on the pre-training phase through estimation of the policy’s uncertainty, and is capable of automatically recovering a “rebalancing” strategy.
+Moreover, AMF-NN considers the reduction in entropy across several tasks: hence, it can in contrast focus on tasks that are harder to learn or that could, in principle, lead to learning progress on other tasks.
+Further empirical evidence for these behaviors is shown in Appendix J.
+Appendix J Single-task performance
+This Section presents a detailed look at the data selection strategies induced by AMF-NN.
+For this purpose, we consider the main experiments in Kitchen and Metaworld outlined in Figure 4, and plot single-task success rates, as well as the amount of demonstrations collected over time.
+For reference, asymptotic single-task success rates on Metaworld and Kitchen converge between 70% and 100%, depending on the task.
+In the case of skewed pre-training (Fig.
+16 and 18), we observe that AMF samples tasks that were not present in the pretraining dataset more often, without having access to any direct information on the pre-training distribution.
+Moreover, even if multiple tasks have the same frequency in the pre-training distribution, AMF will prefer the ones that induce a larger reduction in posterior uncertainty: for instance, in Metaworld, AMF selects Close Faucet more often than Open Faucet, despite the fact that both were similarly demonstrated during pre-training.
+We remark that these task selection strategies arise naturally from our information-based criterion in Equation 1, without any direct information on the pre-training distribution, nor any explicit policy evaluation.
+  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
+Figure 15: Single-task curves for skewed pre-training in Kitchen. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
+  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
+Figure 16: Single-task curves for uniform pre-training in Kitchen. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
+  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
+Figure 17: Single-task curves for skewed pre-training in Metaworld. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
+  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
+Figure 18: Single-task curves for uniform pre-training in Metaworld. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
+Appendix K Analysis of Importance Weights
+Importance weights (as introduced in Equation 3) allow estimating the expert’s occupancy for arbitrary tasks.
+Naturally, the quality of importance weights depends on many factors, including the dimensionality of the trajectory space, and the density with which available data covers it.
+In this section, we report a qualitative evaluation of importance weights for both AMF-GP and AMF-NN (Figures 19 and 20, respectively).
+In both cases, we find that informative weights can be retrieved eventually, given the proper amount of clipping (as described in Appendix M).
+While in early rounds of the algorithm, weights can be inaccurate, leading to a poor estimate of the objective, we observe that the quality of importance sampling weights improves within a handful of rounds.
+Figure 19: Analysis of importance sampling weights for AMF-GP.
+We consider the skewed pre-training setting from Figure 2, and compute importance weights after 1, 5 and 9 rounds.
+We sample four tasks $c_{0:3}$, represented by vertical dashed lines of different colors.
+For each task $c_{i}$, we collect a demonstration $\tau_{i}$ and sweep over $c^{\prime}\in\mathcal{C}$ on the x-axis; we plot $w(\tau_{i},c^{\prime})$ with solid lines.
+We observe that importance weights are uninformative in early parts of training, but converge to more accurate values within a few rounds.
+Figure 20: Analysis of importance sampling weights for AMF-NN.
+We consider the skewed pre-training setting from Figure 4, and compute importance weights after 1, 10 and 19 rounds.
+We visualize weights for both Kitchen (top) and Metaworld (bottom).
+As the task set is discrete, we consider all tasks $(c_{i}\in\mathcal{C})$, and collect one demonstration $\tau_{i}$ for each.
+The entry of each colormap at row $i$ and column $j$ represents $w(\tau_{j},c_{i})$.
+Again, we observe that at the beginning of training importance weights can be inaccurate, particularly for tasks $c_{i}$ that have not been sufficiently demonstrated.
+However, as more data is collected and the policy specializes to each task, the weights converge.
+Appendix L Criterion vs returns
 0.00.10.20.3Pretraining distribution0.00.51.01.51e-4Estimated expected posterior entropyTask space-13.5-13.0-12.5Average return after finetuning
@@
 In this instance, it’s easy to see that the criterion leads to selection of tasks which have not been demonstrated sufficiently, and that will thus lead to greater policy performance.
-Appendix I Implementation Details
-In order to ease reproducibility, we open-source our codebase on the project’s anonymous website.444sites.google.com/active-multitask-finetuning Furthermore, we describe several implementation details in the following sections.
-I.1 Metrics
+Appendix M Implementation Details
+In order to ease reproducibility, we open-source our codebase on the project’s repo.444github.com/marbaga/amf Furthermore, we describe several implementation details in the following sections.
+M.1 Metrics
 All metrics are reported in the form of their mean and the 90% simple bootstrap confidence intervals over $10$ random seeds.
-I.2 GP settings
+M.2 GP settings
 In GP settings (5.1), each expert demonstration involves $5$ steps, is corrupted with Gaussian noise and collected by a scripted policy.
@@
 Multi-task returns are averaged over $20$ episodes per task.
-I.3 Neural network settings
-I.3.1 Environments
-We evaluate AMF-NN across three environments, namely FrankaKitchen, Metaworld and WidowX.
-For the first two, demonstrations are $50$ steps, while for the latter they involve $100$ steps.
+M.3 Neural network settings
+M.3.1 Environments
+We evaluate AMF-NN across four environment suites, namely FrankaKitchen, Metaworld, Robomimic and WidowX.
+For the first two, demonstrations are $\approx 50$ steps, while while for the latter two they involve up to $700$ steps, and $100$ steps respectively.
 In FrankaKitchen, demonstrations are provided by Kumar et al.
@@
 In Metaworld, demostrations are instead collected by the scripted policies provided (Yu et al., 2020).
+In Robomimic, demonstrations are provided by proeficient human demonstrators (Mandlekar et al., 2021).
 Finally, in WidowX successful trajectories are collected by Octo-small (Octo Model Team et al., 2024) itself and filtered according to success labels, in an instance of self-supervised distillation.
 Furthermore, in the case of WidowX, the initial position of the object is not randomized, as we found this to result in very inconsistent performance for the data collection policy.
-In the first two environments, $25$ attempts for each task are evaluated, while the evaluation for WidowX involves $50$ attempts.
-I.3.2 Behavior cloning
-The policy is parameterized via a deterministic MLP with $2$ layers and $256$ units per layer, with layer normalization (Ba, 2016).
-Task conditioning are image embedding extracted by R3M (Nair et al., 2022a).
-Policies are pre-trained for $500$ epochs with batch size of $128$, learning rate of $41.e-$ 4 using the Adam optimizer (Kingma, 2014).
-I.3.3 Importance sampling weights
+In the first two suites, $50$ attempts for each task are evaluated, while evaluation on Robomimic and WidowX involves $25$ attempts.
+M.3.2 Behavior cloning
+The MLP policy has with $2$ layers and $256$ units per layer, with layer normalization (Ba, 2016).
+The Diffusion policy shares architecture and hyperparameters with the original implementation (Chi et al., 2023).
+Task conditioning are word embedding extracted by a sentence transformer (all-MiniLM-L6-v2) (HuggingFace, 2026).
+Policies are pre-trained for $200$ epochs with batch size of $256$, learning rate of $10^{-4}$ using the AdamW optimizer(Loshchilov & Hutter, 2019).
+M.3.3 Importance sampling weights
 In GP setting, importance weights are computed from the Gaussian policy distribution, and log-probabilities are clipped to the range $[-12,0]$.
@@
 We note that this solution cannot be used to evaluate the criterion on yet unobserved tasks, but remains feasible when tasks are finite and few.
-I.3.4 AMF
-Each fine-tuning round involves $3000$ gradient steps, each with a batch size of $128$.
+M.3.4 AMF
+Each fine-tuning round involves $3000$ gradient steps, each with a batch size of $256$.
 We warm-start each algorithm by collecting the first $|\mathcal{C}|$ demonstrations uniformly, as mentioned in Section 4.2.
 In the case of loss-gradient embeddings, we found it to be beneficial to use a separate copy of the policy for task selection, which is not trained on these initial trajectories (which thus can be seen as a small “validation” set).
-As these demonstrations are not selected according to the criterion, to avoid unwanted updates of pre-training weights, they are only added to the training set once the algorithm selects a task belonging to the same partition of the task space, as defined by parameter isolation.
-Appendix J Useful inequalities
-Lemma 8.
+M.3.5 Adaptive Prior
+We found relatively high learning rates for the mixing weight $\alpha$ to work well: all experiments use a learning rate of $1.0$.
+The conservative coefficient $\beta$ was tuned for each suite, and is set to $0.01$ for all but Kitchen, in which it takes the value of $0.001$.
+M.4 Runtime
+Each experimental run for AMF-NN takes at most 5 hours with GPU acceleration.
+In this case, data selection itself requires less than 1 second per round, and can be significantly sped up by reducing the sampling budget.
+AMF-GP experiments can be reproduced within 10 minutes on CPU.
+Appendix N Useful inequalities
+Lemma N.1.
 If $a,b\in(0,M]$ for some $M>0$ and $b\geq a$ then
@@
 $b-a=c\cdot\log\left(\frac{b}{a}\right)>M^{\prime}\cdot\log\left(\frac{b}{a}\right).$
-Lemma 9.
+Lemma N.2.
 Let us consider two spaces $X\in\mathbb{R}^{n}$, $Y\in\mathbb{R}^{m}$, and a conditional distribution $p:X\to\Delta(Y)$ whose support $\text{supp}(p(\cdot|x))$ is bounded by a ball of radius $\epsilon$ for all $x\in X$, that is